### v1

# Keeping in Place After the Storm—Emergency Assistance and Evictions

Abstract
We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided. We find a 16% increase in overall evictions after hurricane natural disaster events that can be decomposed to a 36% increase for those that do not receive FEMA rental assitance and only a 11% for hurricane events that do receive FEMA aid. Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes remain stable and result in a decrease in defaults by 19% in payday loans during hurricane events in locations that do receive FEMA aid. This effect largely reverses in areas that do not receive FEMA aid, where transaction volumes drop by 12% and default rates remain similar relative to the baseline. Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
1 Introduction
The loss of structures and in particularly housing in the wake of natural disasters is often among the notable and lasting images of large natural disasters. Recent catastrophes such as Hurricane Katrina in 2005 and Hurricane Harvey in 2017, continue to attract attention because of the their large-scale visible impacts, notably the destruction of buildings and neighborhoods, the resultant displacement and the post-mortem evaluation of government responses. While the response to such events may yield silver linings in the ensuing learned adaptation, at the micro level, the loss born by individuals can be large and has been documented in various forms in the literature.
Historical instances in which neighborhoods that have been rebuilt do not necessarily resemble their former composition as a result of the immediate loss of housing in such events. The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States. To contrast, in our work, we study instances of assistance in preserving housing access. That is, where federal assistance exists to enable households to maintain their ability to make housing payments when potential disruptions directly related to the storm event affect the ability to make such payments.
Prior to the creation of the Federal Emergency Management Agency under the Carter administration in 1979, many natural disasters that have marked the history of the United States have required federal emergency funding through ad-hoc legislation in the aftermath of their occurrence. One form of disaster relief has been the need for emergency housing in the wake of damage and destruction to housing property. This has further expanded to include rental assistance starting initially under the Stafford Act of 1988 and then further expanded under the Individuals and Households Program in 2002333(Webster, 2022) that aim to recognize disruption in household financial ability due to uninsured or under-insured emergency expenses.
In the same vein, we study the availability of emergency rental assistance through FEMA in the wake of hurricane disaster events and their impact on rental evictions in the state of Florida. We find that notably the rate of eviction is mitigated via the availability of FEMA rental assistance and furthermore we document, in the case of payday loans, that the demand for emergency credit is substituted away from higher cost credit.
1.1 Literature Review
The household financial impacts of natural disasters have been increasingly studied, with significant attention given to household resilience and recovery mechanisms. The intersection of disasters and eviction risks has also garnered attention. (Brennan et al., 2022) explore the heightened vulnerability of renters in disaster-affected areas, showing that disasters exacerbate eviction rates, particularly for vulnerable populations. (Raymond et al., 2022) examine the legal and policy dimensions of eviction laws, emphasizing the gaps in tenant protections in states such as Florida. These findings are particularly relevant to our work, underscoring the cascading effects of disasters on eviction, financial distress, and reliance on high-cost credit.
Furthermore, the consequences of hurricane events specifically has more generally received attention in the literature, (Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances. (Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions. (Del Valle et al., 2024) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster. They identify shifts in financial behavior, including increased reliance on savings and external credit. Similarly, (Collier et al., 2024) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions. Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
The literature on payday loans has highlighted the dual nature of high-cost credit, as both a financial lifeline and a debt trap. We similarly contrast its availability in the context of natural disasters.(Bhutta et al., 2015) describe how payday loans exacerbate financial strain for vulnerable borrowers, while (Morse, 2011) offers a contrasting view, emphasizing their utility during emergencies. (Gathergood et al., 2019) shows that payday loans increase financial distress, especially for low-income borrowers, but regulatory interventions can mitigate harmful effects. In the same vein, (Dobridge, 2016) finds mixed effects, while access to high-cost credit aids short-term liquidity, it negatively affects long-term financial outcomes. Extending this debate, our findings suggest that payday loans act as a substitute for formal disaster relief, especially in regions excluded from FEMA’s coverage. This substitution underscores the dual-edged nature of payday lending in disaster recovery contexts, offering immediate relief but also amplifying financial fragility as we observe a difference in subsequent default rates.
Beyond individual outcomes, the broader economic implications of disasters have been extensively analyzed. (Bernstein et al., 2019) highlight the long-term economic vulnerabilities of disaster-prone areas, particularly the impact of rising sea levels on property prices. The interplay between financial aid and alternative credit sources offers another dimension to understanding recovery dynamics. (Collier et al., 2024) analyze demand for disaster recovery loans, illustrating how credit availability shapes household recovery trajectories. (Lane, 2024) underscores the importance of guaranteed credit programs in building resilience, particularly for lower-income households. (Malmin, 2023) explores the role of federal credit in shaping future wealth trajectories, emphasizing the need for equitable and targeted recovery programs.
Finally, recent studies on disaster aid distribution and its long-term effects highlight systemic inequalities. (Billings et al., 2022) document the inequities in financial aid distribution during Hurricane Harvey, revealing gaps in support for vulnerable populations. (Bufe et al., 2021) explores how financial shocks impact lower-income households, identifying key factors that contribute to financial resilience, such as savings, credit access, and social support networks. (Ratnadiwakara and Venugopal, 2020) finds that flood-prone areas increasingly attract lower-income and less creditworthy populations due to declining property values and housing costs.
Using Florida as a case study, our research extends prior work by integrating multiple granular datasets to examine the dynamics of financial distress through eviction risks and high-cost credit during the recovery following disasters. Our granular spatial and temporal analysis reveals how gaps in disaster aid exacerbate loss of housing via increased evictions, driving households to substitute formal aid with high-cost credit options. This approach reinforces the importance of inclusive disaster recovery programs that address the nuanced needs of vulnerable populations, ultimately potentially mitigating long-term financial harm and fostering resilience.
The rest of the paper is as follows. In section 2, we describe the various data used and our spatial-temporal approach. In section 3 ,we present our main result linking eviction outcomes and access to emergency financial assistance. And then in section 4, we further extend this analysis to understand the interdependence with access to high-cost credit via payday loans.
2 Data and Methodology
2.1 Data
Our data comes from several administrative sources. Our access to evictions data comes from the Evictions Lab at Princeton, which sources their data via county level court records on evictions. Our access to payday loans data comes via a freedom of information request to the state of Florida’s Office of Financial Regulation. We also utilize several federal level datasets. Specifically we retrieve FEMA’s (Federal Emergency Management Agency) Web Disaster Declarations and Housing Assistance Program Data for Renters, as well as publicly available data from NOAA (the National Oceanic and Atmospheric Administration) and the NHC (National Hurricane Center). Finally, estimates of urban and rural population, housing units, and characteristics at the census level come from the American Community Survey.
2.1.1 Eviction Data
Princeton’s Eviction Lab data provides Census-Tract level information at a weekly level, including eviction, filing, and at-risk outcome variables Desmond et al. (2018). Covering the period from 2003 to 2017. Table 1 shows summary statistics. The evict variable describes the number of evictions aggregated at the weekly-census tract level. The filing variable describes the number of court filings for eviction and the atrisk variable adjusts this for the number of individuals living in the household at risk of being evicted.
Table 1: Summary Statistics: Evictions dataset, measures per weekly-census tract.
Statistic evict filing atrisk
25th Perc. 1 1 1
Median 1 1 1
Mean 1.219 1.770 1.676
75th Perc. 1 2 2
Std. Dev. 1.26 1.59 1.49
Observations: 498793 Unique Census Tracts: 3980
2.1.2 Payday Loan Data
Our dataset on payday loans includes loan-level transactions from 2002 to 2018 in the state of Florida, totaling 101 million daily loan-level observations. We are able to observe the loan date, due date, transaction amount, and any related fees. We also observe the date of repayment or whether the loan is outstanding. Furthermore, each loan observation has information on the retailer and borrower’s ZIP code.
2.1.3 Hurricane Forecast Cone
The National Hurricane Center (NHC) provides forecast cones for all hurricane and tropical storm events and includes the following characteristics: the forecast date, hurricane category, and the forecast time of impact. We use these cones to construct a hurricane time-of-impact and path dataset by examining all the cones for all hurricane and tropical storm events that traverse the state of Florida
However, the forecast cones are only available as far back as 2008. For storm events prior to 2008, we utilize NHC’s data on the storm’s realized path to reconstruct the forecast cone for each storm using the same properties as the forecast cones provided by the NHC in the period after 2008.444https://www.nhc.noaa.gov/aboutcone.shtml provides details on how the forecast cone is constructed. Figure 1 shows an example of a forecast cone. The first point represents the actual location of the hurricane, and as we move away from the initial point, the cone covers a larger area with increasing uncertainty, a method of construction chosen by the NHC based on the prediction error over the past five years.
Figure 1: 5-Day Track Forecast Cone
2.1.4 FEMA Data
Since 1979, the Federal Emergency Management Agency (FEMA) has served as the federal government’s lead agency in responding to and facilitating recovery from most significant crises. To provide insight into FEMA’s Housing Assistance Program for renters, the Individual Assistance (IA) reporting team created a dataset beginning with disaster declaration DR1439 (2002). This dataset includes aggregated, non-personally identifiable information such as the number of applications, inspections, the extent of damages, aid provided, and other relevant metrics, broken down by state, county, and ZIP code where registrations were made. FEMA’s Housing Assistance Program for renters, a component of the Individuals and Households Program (IHP), offers critical financial support to those impacted by federally declared disasters. Assistance includes temporary rental aid, Other Needs Assistance (ONA) for uninsured expenses (e.g., personal property or childcare), and Direct Housing Assistance (e.g., transportable housing units or repaired multi-family homes) for cases where rental housing is unavailable.
Before an individual can receive FEMA aid, they must meet specific general eligibility criteria. Applicants are required by FEMA to confirm that the disaster-damaged home is their primary residence that their disaster-related needs are not already covered by another source, such as insurance or other programs. If applicants have insurance, FEMA will require proof of the settlement or a letter explaining why coverage was denied before determining the assistance they are eligible to receive.
We utilize two FEMA datasets: the FEMA Web Disaster Declarations and the Housing Assistance Program Data on Renters. The first dataset includes information on all declared hurricanes and tropical storms affecting Florida, including the disaster name and its corresponding Disaster Number (a sequentially assigned identifier for declared disasters). The Disaster Number is essential for linking to the second dataset, which provides detailed information about housing assistance. We then identify ZIP codes where at least one individual has received rental disaster assistance.
Table 2 presents a list of all major disasters in Florida that were approved for the Individuals and Households Program (IHP). Table 3 provides summary statistics of key variables for the Housing Assistance Program on Renters.
Table 2: Major Disasters and Declaration Types
Declaration Date Disaster Name Declaration Type
2004-08-13 Tropical Storm Bonnie and Hurricane Charley Major Disaster
2004-09-04 Hurricane Frances Major Disaster
2004-09-16 Hurricane Ivan Major Disaster
2004-09-26 Hurricane Jeanne Major Disaster
2005-07-10 Hurricane Dennis Major Disaster
2005-10-24 Hurricane Wilma Major Disaster
2008-08-24 Tropical Storm Fay Major Disaster
2012-07-03 Tropical Storm Debby Major Disaster
2016-09-28 Hurricane Hermine Major Disaster
2017-09-10 Hurricane Irma Major Disaster
Table 3: Summary Statistics by Disaster Name
Disaster Name Registered for FEMA Aid Approved for FEMA Aid Total IHP Amount Disbursed Total Rental Assistance Disbursed Total IHP Amount per Approved Applicant
Tropical Storm Bonnie and Hurricane Charley 50,136 28,945 70,702,271 26,273,487 2,443
Hurricane Frances 114,540 66,784 156,876,161 46,208,593 2,349
Hurricane Ivan 35,961 21,203 52,330,846 11,684,323 2,468
Hurricane Jeanne 89,278 57075 153,826,507 39,083,876 2,695
Hurricane Dennis 9,131 4238 4,018,903 1,006,428 948
Hurricane Wilma 150,501 57,542 92,416,813 23,939,083 1,606
Tropical Storm Fay 5,463 1622 3,395,589 2,098,470 2,093
Tropical Storm Debby 4,455 1,752 4,555,721 2,771,489 2,600
Hurricane Hermine 1,643 585 1,772,365 797,293 3,030
Hurricane Irma 1,420,062 444,735 495,234,951 304,094,846 1,114
2.2 Methodology
We analyze our eviction and payday loan datasets to measure the impact of hurricanes and tropical storms on households. Our focus is on renters, and since our eviction data is at the census tract level, we weighted the outcome variables (evictions, filings, and at-risk individuals) based on the number of renters in each census tract. For the payday loan dataset, we aggregate defaults and number of transactions that we observe at the ZIP code level per day.
We use two models to measure the impact of natural disasters on households. First, we plot event studys to visually analyze the different outcome variables for both the treated and control groups. Then, we used a difference-in-differences model to confirm our results from the event study. Finally, we also then study the existence of federal aid assistance in an interacted difference-in-differences setting to measure the role of FEMA aid.
In the following three sections, we will explain how the treated and control groups were constructed and provide a detailed explanation of how the event study and difference-in-differences models were conducted.
2.2.1 Construction of treatment and control groups
To measure the impact of hurricanes on households, we construct two groups: treatment and control. The treated group consists of ZIP codes (or census tracts) that were within the forecast cone at any point for the entire path of the storm event, from the first prediction to the storm’s dissipation. These areas appear in the forecast cone and are likely to experience some direct or/and indirect effects from the hurricane.
In contrast, our control group includes all ZIP codes (census tracts) that remain completely outside the forecast cone throughout the storm’s progression. This ensures that the behavior and outcomes in these areas were more likely to not be influenced by the forecast or the storm itself, allowing us to isolate the impact of the disaster.
Figures 3 and 3 provide insight into our strategy. They show the evolution of four forecast cones for hurricane Hermine in 2016. The treated group is the green area, whereas the control group is the intersection between all blue areas, as demonstrated in Figure 4.
Figure 2: Forecast cones: distinguish between areas inside and outside the cone
Figure 3: Forecast cones: Evolution of forecast cones during the path of the storm with an example of the treated group (green area)
Figure 4 provides insight into how we constructed the treatment and control groups.
Control group (Blue): This group includes ZIP codes (census tracts) that are entirely outside of any given forecast cone, meaning that these areas were not expected to be affected by a hurricane at any point in time. By focusing on areas not under the forecast at any point, we construct a control group whose behavior and outcomes are assumed to be unaffected directly by storm warnings or the storm itself.
Treated Group (Green): This group includes ZIP codes (census tracts) located within the hurricane forecast cone. However, only areas that remained in the forecast cone up until the closest forecast cone (3 hours prior to impact) are included. That way, we include areas that were eventually either directly in or very close to the realized hurricane path.
In this setup, we compare two groups that likely have very differential impact. We compare the regions that were actually affected (the green area) to the regions (the blue area) that were consistently outside any potential threat of impact. The areas in red were excluded as they were previously in the forecast path but did not eventually remain in the realized path of the storm.
Figure 4: Dynamic 5-Day Forecast Cone
2.2.2 Event Study Framework
To construct the data for the event study, we treat each hurricane and tropical storm as a separate event study. We merge the constructed treated and control group dataset of storm affected geographies with the evictions dataset at the census tract level for six months before and six months after the event. For the treated group, we use the exact date of the storm path for each census tract and merge with the observed week in the eviction dataset. In the control group, since the geographies were constructed to not intersect with the storm forecast, we assign a pseudo-date, in particular, we use the median storm date. Since our evictions dataset is at the weekly level, we calculate the relative event date in weeks and also in months to the storm date observed. We follow the same approach with payday loans, calculating the relative event date at the daily level since our payday loan data is recorded daily.
The estimating equation for our event study model is given by :
$y_{h,i,\tau}=\left(\sum_{j=-m}^{n}\beta_{j}\cdot\large\mathds{1}\{\tau=j\}\cdot D_{h,i}\right)+\alpha_{h,i}+\delta_{h,\tau}+\epsilon_{h,i,\tau}$ (1)
In this equation, $y_{h,i,\tau}$ represents the outcome variable for unit $i$ at relative time $\tau$ for hurricane event $h$. The term $\large\mathds{1}\{\tau=j\}\cdot D_{h,i}$ is an indicator variable that equals 1 if the hurricane event $h$ occurred at relative time to impact $j$ in a treated ZIP code $i$, the unit of observation, and 0 otherwise. The coefficients $\beta_{j}$ capture the dynamic effects of the event over time. For $\tau>0$, these coefficients show the post-event impact, while for $\tau<0$, they provide a measure for pre-event trends. The summation term thus captures the event study terms over all pre- and post-event periods.
The model also includes $\alpha_{h,i}$, which denotes event-by-unit fixed effects to account for event-specific and time-invariant heterogeneity across units, and $\delta_{h,\tau}$, which represents event-by-time fixed effects to control for event-specific time-specific shocks common to all units, ensuring that all the identification comes from within an event. Together, these terms ensure the model effectively isolates the impact of the event while accounting for unobserved heterogeneity.
As in (Doruk Cengiz and Zipperer, 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously. This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
To properly isolate both treated and control groups and to prevent overlap, we only consider a ZIP code as part of the treated group when it is the first event affecting each ZIP code during the event window. Essentially, we exclude ZIP codes that experience treatment of any other storm events during the construction of the event window of the control group. Similarly, for the control group, we only consider the never-treated during the event window.
In our setup, we conduct multiple event studies focusing on evictions, filings, households at risk, transaction volumes, and default rates for both treated and control groups.
For the evictions data, we used event-specific ZIP codes as the geographic fixed effect. Considering the eviction process spans potentially several months from initial notice, to filing, to eviction, we measured the effect spanning six months. For the payday loans data, we aggregated daily observations at the weekly level to account for within week seasonality, we used event-specific ZIP codes as the geographic fixed effect and event-specific relative weeks as the time fixed effects.
2.2.3 Difference-in-Differences Framework
We also present the event-study results in a difference-in-difference analysis comparing the treated and control groups in pre and post-event following each event study. The model is as follows:
$y_{h,i,\tau}=\beta\cdot\large\mathds{1}\{\tau\geq 0\}\cdot D_{h,i}+\alpha_{h,i}+\delta_{h,\tau}+\epsilon_{h,i,\tau}$ (2)
where $y_{h,i,\tau}$ denotes the dependent variable, representing (evictions, filings, transaction volume, defaults) for observation of hurricane event $h$, unit $i$, at relative time $\tau$. $D_{h,i}$ is an indicator variable for the treatment group, equals one for the treated ZIP code and zero otherwise for hurricane event $h$. The indicator variable $\large\mathds{1}\{\tau\geq 0\}$ represents the post-treatment period. By constructing the panel based on relative time, we ensure a framework in which events are aligned to occur simultaneously. And $\alpha_{h,i},\delta_{h,\tau}$ denotes the event specific fixed effects for unit and time (ZIP code and week/month), respectively.
Finally, to shed light on the role of FEMA aid, we conduct an interacted difference-in-differences comparison for hurricane events that are declared FEMA and resulted in the availability of rental assistance (or were declared FEMA but did not result in any rental assistance). For this approach, the control group is the same as the event based matched control group in the earlier difference-in-difference; in contrast, we divide the treated group into two groups. The non-FEMA treated group contains ZIP codes that were impacted by hurricanes or tropical storms but didn’t receive FEMA aid, and the treated FEMA group is ZIP codes impacted by hurricanes and tropical storms that received FEMA aid.
In the next section, we present the findings from the eviction and payday loan analysis, focusing on the impact of hurricanes and FEMA aid on eviction, transactions and defaults.
3 Results
3.1 Evictions
3.1.1 Hurricanes
Using our evictions dataset we find that ZIP codes that experienced hurricanes saw a monthly average increase of evictions of 0.525 per month per 1000 renters [^5] We normalize the raw number of evictions and filings using census data on the number of renters present in a ZIP code to account for heterogeneity in the percentage of renters and in the population per ZIP code. Table 4. This represents a monthly increase of 16% evictions per month compared to the baseline. The event study in Figure 5 shows that this effect persists up to 6 months after the initial hurricane. The event study also shows a gradual increase in the initial two months up to a peak around months 3-5, reflecting the nature of the eviction process that may lag from the initial missed rental payment date, to the eviction filing, to the finally court approved eviction order.
We also evaluate the change in the number of filings, in which we find an increase of 0.928 in the number of filings per month per 1000 renters, in the 6 month that follow the hurricane event. This represents a 28.7% increase relative to the baseline rate of filings per month.
Figure 5: The figure provides the estimated coefficients from an event study examining the effect of hurricane exposure on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. Coefficients for Treated and Post are not reported, as they are absorbed by fixed effects. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. The coefficient of 0.525 for evictions represents an approximate 16.2% increase relative to the treated group’s pre-event average. Similarly, the 0.928 increase in filings corresponds to about a 17.9% rise from the treated group’s pre-event average. Standard errors clustered at the ZIP code level are reported in square brackets. ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
[0.172] [0.235]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 15,641 15,641
Table 5: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
Variable Control Group Treated Group
Eviction Rate 3.665 3.238
Filing Rate 5.592 5.182
We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes. Our results show a much larger bifurcation in the resultant filings and evictions in the period following the hurricane event. Both the event-study and difference-in-differences regression results illuminate the increased rate of filings and evictions for hurricane events where FEMA assistance is not made available. In particular, we find an increased monthly average of 1.115 and 1.543, for evictions and filings per month, respectively Table 6. These represent a 35.7% and 29.8% increase, respectively, for hurricanes that did not receive FEMA aid. Comparatively, the availability of FEMA rental assistance resulted in a 11.2% increase over the baseline in evictions. The absence of FEMA assistance resulted in three times the increase in the rate of evictions and two times the increase in the rate of filings. This also highlights that there was not only a greater increase in the number of evictions but a comparatively greater increase in the rate of successful evictions relative to filings, given the greater increase of evictions to filings.
Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction rates at the ZIP code level. The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the monthly eviction rate. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level. The estimates suggest a marked post-event increase in eviction rates in non-recipient ZIP codes, whereas areas receiving aid experienced relatively muted changes, indicating a potential mitigating role of disaster assistance.
Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly filing rates at the ZIP code level. The event window covers five months prior to and six months following the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays results for ZIP codes that received FEMA aid, while Panel (b) shows results for ZIP codes that did not receive FEMA aid. The dependent variable is the monthly filing rate. Vertical lines indicate 95% confidence intervals, calculated using standard errors clustered at the ZIP code level. The findings reveal a substantial post-event increase in filing rates in non-recipient areas, whereas the increase is more muted in aid-recipient ZIP codes—suggesting that FEMA assistance may play a mitigating role in the aftermath of natural disasters.
Table 6: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, estimated separately for FEMA and non-FEMA designated areas. The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes. For FEMA areas, the coefficient of 0.363 for evictions represents an approximate 10.6% increase relative to the treated group’s pre-event average (3.427), while the 0.769 increase in filings corresponds to about a 14.3% rise from the treated group’s baseline (5.392). In non-FEMA areas, the 1.155 increase in evictions indicates a roughly 50.1% rise from the treated pre-event average (2.307), and the 1.543 increase in filings represents a 37.2% increase from the baseline (4.143).
Dependent Variable Eviction (FEMA) Filing (FEMA) Eviction (No-FEMA) Filing (No-FEMA)
Period Monthly Monthly Monthly Monthly
Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.363 0.769∗∗∗ 1.155∗∗∗ 1.543∗∗∗
[0.188] [0.251] [0.353] [0.510]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 11,097 11,097 4,544 4,544
Table 7: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
Variable Control Group Treated Group
FEMA No FEMA FEMA No FEMA
Eviction Rate 3.578 3.869 3.427 2.307
Filing Rate 5.298 6.273 5.392 4.143
Table 8: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level. The specification includes an interaction between Treated, Post, and a FEMA designation indicator, allowing the treatment effect to vary by FEMA status. The estimation window spans 5 months before to 6 months after the hurricane event. The coefficient on Treated×$\times$×Post indicates the average treatment effect in non-FEMA areas: a 1.155 increase in monthly evictions (approximately 50.1% relative to the treated pre-event mean of 2.307) and a 1.543 increase in filings (a 37.2% increase from the baseline of 4.143). The negative Treated×$\times$×Post×$\times$×FEMA interaction suggests that the treatment effect was significantly smaller in FEMA-designated areas: 0.792 fewer evictions and 0.775 fewer filings, relative to non-FEMA treated ZIP codes.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 1.155∗∗∗ 1.543∗∗∗
[0.353] [0.510]
Treated×$\times$×Post×$\times$×FEMA -0.792∗∗ -0.775
[0.388] [0.548]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 15,641 15,641
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
3.1.2 Hurricanes and Tropical Storms
Table 9: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
Variable Control Group Treated Group
Eviction Rate 3.937 3.113
Filing Rate 5.808 5.033
We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included. Although, the comparative magnitudes decreases and and the difference due to availability of FEMA assistance also is not statistically different than zero. We show that this is largely due to tropical storms seeing far lower post-storm eviction rates. Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events. Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the absence of FEMA assistance.
Figure 8 and Table 10 shows the baseline result even when including the greater number of tropical storm events still holds (note the three times increase in sample size). The decrease is largely driven by the inclusion of tropical storms, noting the treatment post-period coefficient drop from 0.371 to 0.128 for evictions when only tropical storms are included, although the decrease itself is not statistically significant.
Figure 8: The figure provides the estimated coefficients from an event study examining the effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level. The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level. The results indicate a notable post-event increase in both eviction and filing rates, highlighting the disruptive impact of storm exposure on evictions.
Figure 9: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on monthly eviction-related outcomes at the ZIP code level. The analysis spans a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period. Panel (a) reports estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level. The results suggest no statistically significant change in either eviction or filing rates following a tropical storm event.
Table 10: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storms and another excluding hurricane events. Coefficients for Treated and Post are not reported, as they are absorbed by event-specific ZIP code and time fixed effects. The estimation window spans from 5 months before to 6 months after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.371 evictions and 0.633 filings in treated ZIP codes. These correspond to approximately 11.9% and 12.6% increases relative to the treated group’s pre-event averages of 3.113 and 5.033, respectively. When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
Dependent Variable Eviction Filing Eviction Filing
Period Monthly Monthly Monthly Monthly
Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.371∗∗∗ 0.633∗∗∗ 0.128 0.171
[0.138] [0.186] [0.218] [0.283]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Hurricanes Y Y N N
Observations 48,076 48,076 32,435 32,435
Table 11: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event. The coefficient of 0.525 for evictions and 0.928 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 16.9% and 18.4% increases relative to the pre-event averages, respectively. The interaction term Treated×$\times$×Post×$\times$×Storm represents the differential impact of storms. For evictions, the coefficient is -0.398, not statistically significant, while for filings, the coefficient is -0.757∗∗, indicating a statistically significant reduction in filings in treated ZIP codes exposed to tropical storms.
Dependent Variable Eviction Filing
Period Monthly Monthly
Window [-5 ; 6] [-5 ; 6]
Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
[0.172] [0.235]
Treated×$\times$×Post×$\times$×Storm -0.398 -0.757∗∗
[0.271] [0.360]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 48,076 48,076
3.2 Payday Loans
3.2.1 Hurricanes
We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance. We find that when FEMA assistance is made available, transactions volumes remain stable while default rates decrease comparatively, reflecting both a reduction in supply of credit when FEMA is not available but also an increase in precarity in making repayments. In particular, Figure 11 and Figure 12 shows that around weeks 6-10 post treatment we observe a spike in defaults and reduction in transaction volumes (possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 15-day loans). This also coincides around when evictions peak for non-FEMA receiving hurricanes in the previous section.
Specifically, in Table 13 we show that the volume of transactions reduces by 8.364 per week and defaults decrease overall by 1.458 per week on average per ZIP code. This reflects a 3.8% decrease in transactions and a 14% decrease in defaults per ZIP code per week relative to the baseline period. Comparatively, when FEMA rental assistance is not available available, transactions volumes decreased by 27.326 per week, a decline of 12.3%. While when FEMA assistance is available, defaults reduced by 1.941 per ZIP code per week relative to the baseline period, a decrease of 18.6% per week per ZIP code relative to the baseline period. And as Table 16 show, on the shorter interval of 7 weeks post hurricane event, this difference was even larger, an overall difference of 29.7% in weekly defaults between hurricane events in which FEMA rental assistance was and was not available.
Table 12: Baseline average weekly level per ZIP code in the pre-event period.
Variable Control Group Treated Group
Transaction Volume 228.833 221.300
Default 8.714 10.418
Figure 10: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the transaction volume. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the default rates. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Table 13: Difference-in-Differences results for transaction volume and default rate outcomes at the ZIP code level for payday loan data. The analysis evaluates the impact of hurricane event across treated and control groups over two different weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -4.857∗∗ for transaction volume in the [-5 ; 7] window suggests a significant decrease in transaction volume in the treated ZIP codes, corresponding to approximately 2.2% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -1.088∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 10.4% lower default rates in treated ZIP codes. The longer window, [-5 ; 12], shows even stronger effects with a coefficient of -8.364∗∗∗ for transaction volume, suggesting a 3.8% reduction in transactions, and -1.458∗∗∗ for default rates, corresponding to a 14% reduction in defaults.
Weekly Window [-5 ; 7] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -4.857∗∗ -1.088∗∗∗ -8.364∗∗∗ -1.458∗∗∗
[2.040] [0.403] [2.121] [0.416]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 25856 25856 35912 35912
Table 14: Baseline average weekly level per ZIP code in the pre-event period.
Variable Control Group Treated Group
FEMA No FEMA FEMA No FEMA
Transaction Volume 209.930 283.549 226.306 202.092
Default 8.286 9.954 11.117 7.738
Table 15: Difference-in-Differences analysis of transaction volume and default rate outcomes at the ZIP code level, disaggregated by FEMA assistance. The analysis compares the impact of the hurricane event on ZIP codes with and without FEMA assistance over two weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]). For the [-5 ; 7] window, ZIP codes with FEMA assistance experienced a small reduction in transaction volume (-2.998), while those without FEMA assistance saw a much larger decrease (-10.675∗∗∗). For default rates, FEMA-assisted ZIP codes experienced a slight reduction (-1.837∗∗∗), while those without FEMA saw an increase (1.253∗∗∗). For the longer [-5 ; 12] window, FEMA-assisted ZIP codes experienced a smaller reduction in transaction volume (-2.356), while non-FEMA-assisted ZIP codes saw a substantial decrease (-27.237∗∗∗). Default rates followed a similar pattern, with a decrease in FEMA-assisted ZIP codes (-1.941∗∗∗) and a small increase in non-FEMA ZIP codes (0.056).
Weekly Window [-5 ; 7]
Dependent Variable Transaction Default
FEMA No FEMA FEMA No FEMA
Treated×$\times$×Post -2.998 -10.675∗∗∗ -1.837∗∗∗ 1.253∗∗∗
[2.750] [1.509] [0.548] [0.327]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 19517 6339 19517 6339
Weekly Window [-5 ; 12]
Dependent Variable Transaction Default
FEMA No FEMA FEMA No FEMA
Treated×$\times$×Post -2.356 -27.237∗∗∗ -1.941∗∗∗ 0.056
[2.648] [2.878] [0.531] [0.827]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 27140 8772 27140 8772
Table 16: Triple Difference-in-Differences (DDD) analysis of transaction volume and default rate outcomes at the ZIP code level, examining the interaction between FEMA assistance and post-event outcomes. The table presents results for two weekly windows: [-5; 7] and [-5; 12]. For both windows, the treatment effect (Treated×$\times$×Post) is negative and statistically significant for transaction volume, indicating a post-event decline in economic activity. Conversely, the treatment effect is positive for default rates, suggesting an increase in loan defaults following the event. The interaction term (Treated×$\times$×Post×$\times$×FEMA) indicates that FEMA-assisted ZIP codes experienced a significant decrease in default rates in both windows, highlighting the mitigating effect of federal disaster assistance.
Weekly Window [-5 ; 7] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -10.675∗∗∗ 1.253∗∗∗ -27.237∗∗∗ 0.056
[1.509] [0.327] [2.877] [0.258]
Treated×$\times$×Post×$\times$×FEMA 7.677∗∗ -3.090∗∗∗ 24.881∗∗ -1.997∗∗∗
[3.088] [0.664] [4.142] [0.607]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 25856 25856 35912 35912
3.2.2 Hurricanes and Tropical Storms Combined
Table 17: Baseline average weekly level per ZIP code in the pre-event period for hurricanes and tropical storms combined
Variable Control Group Treated Group
Transaction Volume 267.802 256.220
Default 8.681 9.537
Table 18: Baseline average weekly level per ZIP code in the pre-event period for tropical storms
Variable Control Group Treated Group
Transaction Volume 286.228 291.139
Default 8.666 8.655
Adding lower intensity tropical storms once again does not overturn the results in comparing events that include FEMA assistance. We continue to observe a decrease in transaction volume and a comparative decrease in default rates for events that saw FEMA assistance. Hence, allowing us to control for possible additional temporal events and for differences possibly due to intensity of non-FEMA hurricane events.
Table 19 and Table 20 show that overall transaction volumes decrease by 7.468 weekly transactions per ZIP code and that defaults decrease by 0.982 defaults per week on average relative to a statistically non-significant change in defaults when only including tropical storm events.
Most of the change in decreases in transaction volumes and in preservation of default rates is not attributed to non-FEMA events being possibly due to lower intensity storms relative to FEMA declared hurricane events as tropical storms show similar magnitudes of changes in transaction volumes and non-decrease in default rates to non-FEMA hurricanes as in Table 15.
Figure 13: The figure provides the estimated coefficients from an event study examining the impact of storm (hurricane and tropical storm) exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Figure 14: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates. Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
Table 19: Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis includes two specifications: one with all storm events (hurricanes and tropical storms), and another excluding hurricane events. The estimation window spans from 5 weeks before to 12 weeks after the event. In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant decline of 7.468 transactions and 0.932 defaults in treated ZIP codes. These correspond to approximately 2.9% and 9.8% decreases relative to the treated group’s pre-event averages of 256.220 transactions and 9.537 defaults, respectively. When hurricanes are excluded, the estimated effects are smaller: a 6.345 decrease in transactions (2.2%) and a non-significant 0.273 decrease in defaults (3.2%).
Weekly Window [-5 ; 12] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -7.468∗∗∗ -0.932∗∗∗ -6.345∗∗∗ -0.273
[1.316] [0.234] [1.847] [0.200]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Hurricanes Y Y N N
Observations 103,291 103,291 67,379 67,379
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
Table 20: Triple Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms. The analysis investigates heterogeneity in the treatment effect by interacting the treatment indicator with a binary storm type indicator (1 = tropical storm, 0 = hurricane). The estimation window spans from 5 weeks before to 12 weeks after the event. The coefficient on Treated×$\times$×Post indicates a statistically significant decrease of 8.364 transactions and 1.458 defaults in treated ZIP codes overall. The interaction term Treated×$\times$×Post×$\times$×Storm captures differential effects by storm type: default rates rise after tropical storms relative to hurricanes, potentially reflecting the role of FEMA assistance, which is less commonly triggered by tropical storms.
Weekly Window [-5 ; 12]
Dependent Variable Transaction Default
Treated×$\times$×Post -8.364∗∗∗ -1.458∗∗∗
[2.040] [0.403]
Treated×$\times$×Post×$\times$×Storm 2.019 1.185∗∗
[2.907] [0.463]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 103,291 103,291
${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
4 Conclusion
Using granular data on evictions and loan-level data on payday loans, we provide evidence of disaster induced increase in evictions and reliance on high cost credit in the state of Florida following major storms and hurricanes. We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit. Overall, this suggests that major individual level household impacts such as access to housing are sensitive to the availability of emergency liquidity, supporting the need for such programs and the need to understand the availability of other forms of disaster insurance in a world where many locations will be subject to an increasing propensity of climate shock events.

### v2

@@
 We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided.
-We find a 16% increase in overall evictions after hurricane natural disaster events that can be decomposed to a 36% increase for those that do not receive FEMA rental assitance and only a 11% for hurricane events that do receive FEMA aid.
-Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes remain stable and result in a decrease in defaults by 19% in payday loans during hurricane events in locations that do receive FEMA aid.
-This effect largely reverses in areas that do not receive FEMA aid, where transaction volumes drop by 12% and default rates remain similar relative to the baseline.
+We find an approximate 16.5% increase in overall evictions after hurricane natural disaster events that increases to 19.7% when excluding ZIP codes that receive FEMA rental assistance.
+Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes and defaults decrease during hurricane events in locations that do receive FEMA aid.
+This effect largely reverses in areas that do not receive FEMA aid, where the magnitude of transaction volumes drop by less and default rates remain similar relative to the baseline.
 Overall, this suggests that the availability of emergency liquidity during natural disaster events is indeed a binding constraint with real household financial consequences, in particular through our documented channel of evictions and in usage of high-cost credit.
 1 Introduction
-The loss of structures and in particularly housing in the wake of natural disasters is often among the notable and lasting images of large natural disasters.
-Recent catastrophes such as Hurricane Katrina in 2005 and Hurricane Harvey in 2017, continue to attract attention because of the their large-scale visible impacts, notably the destruction of buildings and neighborhoods, the resultant displacement and the post-mortem evaluation of government responses.
-While the response to such events may yield silver linings in the ensuing learned adaptation, at the micro level, the loss born by individuals can be large and has been documented in various forms in the literature.
-Historical instances in which neighborhoods that have been rebuilt do not necessarily resemble their former composition as a result of the immediate loss of housing in such events.
+The loss of structures and in particularly housing is often among the notable and lasting images of in the wake of natural disasters.
+The resulting financial loss and displacement highlight critical questions over both the individual and government role in pre-disaster mitigation and post-disaster response, in particular through their shortcomings.
+At the micro level, the loss born by households can be large and has been documented in various forms in the literature.
+Historically, the destruction of housing in such events has often led to rebuilt neighborhoods that differ significantly from their original form and community composition.
 The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States.
 To contrast, in our work, we study instances of assistance in preserving housing access.
-That is, where federal assistance exists to enable households to maintain their ability to make housing payments when potential disruptions directly related to the storm event affect the ability to make such payments.
+That is, we examine the existence of federal assistance households to maintain their ability to make housing payments due to potential disruptions and damages directly related to the natural disaster.
 Prior to the creation of the Federal Emergency Management Agency under the Carter administration in 1979, many natural disasters that have marked the history of the United States have required federal emergency funding through ad-hoc legislation in the aftermath of their occurrence.
@@
 These findings are particularly relevant to our work, underscoring the cascading effects of disasters on eviction, financial distress, and reliance on high-cost credit.
-Furthermore, the consequences of hurricane events specifically has more generally received attention in the literature, (Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances.
+Furthermore, the consequences of hurricane events has notably received attention in the literature.
+(Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances.
 (Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions.
@@
 2005-07-10 Hurricane Dennis Major Disaster
+2005-08-28 Hurricane Katrina Major Disaster
 2005-10-24 Hurricane Wilma Major Disaster
@@
 Table 3: Summary Statistics by Disaster Name
-Disaster Name Registered for FEMA Aid Approved for FEMA Aid Total IHP Amount Disbursed Total Rental Assistance Disbursed Total IHP Amount per Approved Applicant
-Tropical Storm Bonnie and Hurricane Charley 50,136 28,945 70,702,271 26,273,487 2,443
-Hurricane Frances 114,540 66,784 156,876,161 46,208,593 2,349
-Hurricane Ivan 35,961 21,203 52,330,846 11,684,323 2,468
-Hurricane Jeanne 89,278 57075 153,826,507 39,083,876 2,695
-Hurricane Dennis 9,131 4238 4,018,903 1,006,428 948
-Hurricane Wilma 150,501 57,542 92,416,813 23,939,083 1,606
-Tropical Storm Fay 5,463 1622 3,395,589 2,098,470 2,093
-Tropical Storm Debby 4,455 1,752 4,555,721 2,771,489 2,600
-Hurricane Hermine 1,643 585 1,772,365 797,293 3,030
-Hurricane Irma 1,420,062 444,735 495,234,951 304,094,846 1,114
+Disaster Name
+Registered for
+FEMA Aid
+Approved for
+FEMA Aid
+Total Rental
+Assistance Disbursed (\$)
+Average Rental Assistance
+per Recipient (\$)
+Tropical Storm Bonnie and
+Hurricane Charley
+50,136 28,945 26,273,487 912
+Hurricane Frances 114,540 66,784 46,208,593 695
+Hurricane Ivan 35,961 21,203 11,684,323 554
+Hurricane Jeanne 89,278 57075 39,083,876 687
+Hurricane Dennis 9,131 4238 1,006,428 242
+Hurricane Wilma 150,501 57,542 23,939,083 418
+Tropical Storm Fay 5,463 1,622 2,098,470 1,342
+Tropical Storm Debby 4,455 1,752 2,771,489 1,638
+Hurricane Hermine 1,643 585 797,293 1,442
+Hurricane Irma 1,420,062 444,735 304,094,846 685
 2.2 Methodology
@@
 Together, these terms ensure the model effectively isolates the impact of the event while accounting for unobserved heterogeneity.
-As in (Doruk Cengiz and Zipperer, 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously.
+As in (Cengiz et al., 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously.
 This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
@@
 And $\alpha_{h,i},\delta_{h,\tau}$ denotes the event specific fixed effects for unit and time (ZIP code and week/month), respectively.
-Finally, to shed light on the role of FEMA aid, we conduct an interacted difference-in-differences comparison for hurricane events that are declared FEMA and resulted in the availability of rental assistance (or were declared FEMA but did not result in any rental assistance).
-For this approach, the control group is the same as the event based matched control group in the earlier difference-in-difference; in contrast, we divide the treated group into two groups.
-The non-FEMA treated group contains ZIP codes that were impacted by hurricanes or tropical storms but didn’t receive FEMA aid, and the treated FEMA group is ZIP codes impacted by hurricanes and tropical storms that received FEMA aid.
+Finally, to shed light on the role of FEMA aid, we conduct an interacted event study comparison on the treated ZIP codes for hurricane events and resulted in ZIP codes having the availability of rental assistance.
+In addition we also conducted a difference-in-differences analysis while excluding treated units that received any FEMA assistance, therefore only on the non-FEMA receiving ZIP codes.
+For the difference-in-difference approach, the control group is the same as the event based matched control group in the earlier difference-in-difference; in contrast, for the FEMA event-study on the treated, we separated ZIP codes into those that received FEMA assistance and those do not within the same hurricane event.
 In the next section, we present the findings from the eviction and payday loan analysis, focusing on the impact of hurricanes and FEMA aid on eviction, transactions and defaults.
@@
 Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level.
-Coefficients for Treated and Post are not reported, as they are absorbed by fixed effects.
 The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
-The coefficient of 0.525 for evictions represents an approximate 16.2% increase relative to the treated group’s pre-event average.
-Similarly, the 0.928 increase in filings corresponds to about a 17.9% rise from the treated group’s pre-event average.
+the coefficient of 0.331 for evictions represents an approximate 16.5% increase relative to the treated group’s pre-event average.
+Similarly, the 0.428 increase in filings corresponds to about a 13.3% rise from the treated group’s pre-event average.
 Standard errors clustered at the ZIP code level are reported in square brackets.
@@
 Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
-[0.172] [0.235]
+Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
+[0.089] [0.110]
 Event-specific ZIP Code FE Y Y
 Event-specific Time FE Y Y
-Observations 15,641 15,641
+Observations 18501 18501
 Table 5: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
 Variable Control Group Treated Group
-Eviction Rate 3.665 3.238
-Filing Rate 5.592 5.182
+Eviction Rate 2.403 2.002
+Filing Rate 3.601 3.203
 We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes.
-Our results show a much larger bifurcation in the resultant filings and evictions in the period following the hurricane event.
-Both the event-study and difference-in-differences regression results illuminate the increased rate of filings and evictions for hurricane events where FEMA assistance is not made available.
-In particular, we find an increased monthly average of 1.115 and 1.543, for evictions and filings per month, respectively Table 6.
-These represent a 35.7% and 29.8% increase, respectively, for hurricanes that did not receive FEMA aid.
-Comparatively, the availability of FEMA rental assistance resulted in a 11.2% increase over the baseline in evictions.
-The absence of FEMA assistance resulted in three times the increase in the rate of evictions and two times the increase in the rate of filings.
-This also highlights that there was not only a greater increase in the number of evictions but a comparatively greater increase in the rate of successful evictions relative to filings, given the greater increase of evictions to filings.
-Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction rates at the ZIP code level.
+Our results indicate a possible bifurcation in the resultant filings and evictions in the period following the hurricane event.
+Both the event-study and difference-in-differences regression results illuminate the decrease rate of filings and evictions for hurricane events where FEMA assistance is made available.
+In particular, we firstly find a decrease in the monthly average of 0.418 and 0.445, for evictions and filings per month, respectively Table 6 when restricting to the event study just within treated units.
+These represent a 20.9% and 13.9% decrease, respectively, for hurricanes that did not receive FEMA aid.
+Although both results are not statistically significant, we also show a larger estimate for evictions and filings when excluding ZIP codes that included FEMA assistance in the difference-in-differences results relative to the baseline.
+Comparatively, evictions increased by a statistically significant 0.392 when excluding FEMA receiving ZIP codes relative to the control group, a larger coefficient than 0.331 result when including all ZIP codes.
+Similarly, filings increased by a statistically significant 0.502 when excluding FEMA assistance, larger than the 0.428 coefficient when including all ZIP codes.
+Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level on the treated.
 The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period.
-Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid.
-The dependent variable is the monthly eviction rate.
-Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
-The estimates suggest a marked post-event increase in eviction rates in non-recipient ZIP codes, whereas areas receiving aid experienced relatively muted changes, indicating a potential mitigating role of disaster assistance.
-Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly filing rates at the ZIP code level.
-The event window covers five months prior to and six months following the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period.
-Panel (a) displays results for ZIP codes that received FEMA aid, while Panel (b) shows results for ZIP codes that did not receive FEMA aid.
-The dependent variable is the monthly filing rate.
-Vertical lines indicate 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-The findings reveal a substantial post-event increase in filing rates in non-recipient areas, whereas the increase is more muted in aid-recipient ZIP codes—suggesting that FEMA assistance may play a mitigating role in the aftermath of natural disasters.
-Table 6: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, estimated separately for FEMA and non-FEMA designated areas.
-The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
-For FEMA areas, the coefficient of 0.363 for evictions represents an approximate 10.6% increase relative to the treated group’s pre-event average (3.427), while the 0.769 increase in filings corresponds to about a 14.3% rise from the treated group’s baseline (5.392).
-In non-FEMA areas, the 1.155 increase in evictions indicates a roughly 50.1% rise from the treated pre-event average (2.307), and the 1.543 increase in filings represents a 37.2% increase from the baseline (4.143).
-Dependent Variable Eviction (FEMA) Filing (FEMA) Eviction (No-FEMA) Filing (No-FEMA)
-Period Monthly Monthly Monthly Monthly
-Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.363 0.769∗∗∗ 1.155∗∗∗ 1.543∗∗∗
-[0.188] [0.251] [0.353] [0.510]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Observations 11,097 11,097 4,544 4,544
-Table 7: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
-Variable Control Group Treated Group
-FEMA No FEMA FEMA No FEMA
-Eviction Rate 3.578 3.869 3.427 2.307
-Filing Rate 5.298 6.273 5.392 4.143
-Table 8: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level.
-The specification includes an interaction between Treated, Post, and a FEMA designation indicator, allowing the treatment effect to vary by FEMA status.
-The estimation window spans 5 months before to 6 months after the hurricane event.
-The coefficient on Treated×$\times$×Post indicates the average treatment effect in non-FEMA areas: a 1.155 increase in monthly evictions (approximately 50.1% relative to the treated pre-event mean of 2.307) and a 1.543 increase in filings (a 37.2% increase from the baseline of 4.143).
-The negative Treated×$\times$×Post×$\times$×FEMA interaction suggests that the treatment effect was significantly smaller in FEMA-designated areas: 0.792 fewer evictions and 0.775 fewer filings, relative to non-FEMA treated ZIP codes.
+Panel (a) presents the interacted monthly eviction estimates for ZIP codes that received FEMA aid versus those that did not for the treated group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that received FEMA aid.
+Standard errors are clustered at the ZIP code level.
+Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level with the entire set of control ZIP codes but excluding the treated ZIP codes that do not receive FEMA assistance.
+The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period.
+Panel (a) presents the interacted monthly eviction estimates for ZIP codes that do not receive FEMA assistance versus the control group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that do not receive FEMA assistance versus the control group.
+Standard errors are clustered at the ZIP code level.
 Dependent Variable Eviction Filing
@@
 Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 1.155∗∗∗ 1.543∗∗∗
-[0.353] [0.510]
-Treated×$\times$×Post×$\times$×FEMA -0.792∗∗ -0.775
-[0.388] [0.548]
+FEMA×$\times$×Post -0.418 -0.445
+[0.291] [0.303]
 Event-specific ZIP Code FE Y Y
 Event-specific Time FE Y Y
-Observations 15,641 15,641
-${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
+Observations 2729 2729
+Table 6: Results for monthly eviction and filing outcomes at the ZIP code level on the treated group only, separated into ZIP codes that received FEMA assistance and those did not.
+this provides within storm event estimate comparisons for all ZIP codes that were within the storm path.
+The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
+For FEMA receiving ZIP codes, the coefficient of -0.418, for evictions, while not statistically significant, represents an approximate 20.9% decrease relative to the treated group’s pre-event average (2.002), while the -0.445 decrease in filings corresponds to about a 13.9% decrease from the treated group’s baseline (3.203).
+Dependent Variable Eviction Filing
+Period Monthly Monthly
+Window [-5 ; 6] [-5 ; 6]
+Treated×$\times$×Post 0.392∗∗∗ 0.502∗∗∗
+[0.092] [0.116]
+Event-specific ZIP Code FE Y Y
+Event-specific Time FE Y Y
+Observations 16647 16647
+Table 7: Difference-in-difference results for monthly eviction and filing outcomes at the ZIP code level while excluding treated ZIP codes that received FEMA assistance.
+This provides a comparison with control group ZIP codes that for the same time period were not in the storm path.
+The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
+For non-FEMA receiving ZIP codes, the statistically significant coefficient of 0.392, for evictions, is larger than the 0.331 coefficient for all treated ZIP codes, represents an approximate 19.6% increase relative to the treated group’s pre-event average (2.002), while the 0.502 increase in filings, larger than the 0.428 coefficient for all treated ZIP codes, corresponds to about a 15.6% increase from the treated group’s baseline (3.203).
 3.1.2 Hurricanes and Tropical Storms
@@
 Variable Control Group Treated Group
-Eviction Rate 3.937 3.113
-Filing Rate 5.808 5.033
+Eviction Rate 2.664 2.423
+Filing Rate 3.950 3.829
 We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included.
-Although, the comparative magnitudes decreases and and the difference due to availability of FEMA assistance also is not statistically different than zero.
-We show that this is largely due to tropical storms seeing far lower post-storm eviction rates.
+Although, the comparative magnitudes decreases, we show that this is largely due to tropical storms seeing far lower post-storm eviction rates.
 Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events.
-Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the absence of FEMA assistance.
+Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance.
 Figure 8 and Table 9 shows the baseline result even when including the greater number of tropical storm events still holds (note the three times increase in sample size).
@@
 Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
-The results indicate a notable post-event increase in both eviction and filing rates, highlighting the disruptive impact of storm exposure on evictions.
 Figure 9: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on monthly eviction-related outcomes at the ZIP code level.
@@
 The analysis includes two specifications: one with all storms and another excluding hurricane events.
-Coefficients for Treated and Post are not reported, as they are absorbed by event-specific ZIP code and time fixed effects.
 The estimation window spans from 5 months before to 6 months after the event.
-In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.371 evictions and 0.633 filings in treated ZIP codes.
-These correspond to approximately 11.9% and 12.6% increases relative to the treated group’s pre-event averages of 3.113 and 5.033, respectively.
+In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.209 evictions and 0.336 filings in treated ZIP codes.
+These correspond to approximately 8.6% and 8.7% increases relative to the treated group’s pre-event averages of 2.423 and 3.829, respectively.
 When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
@@
 Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.371∗∗∗ 0.633∗∗∗ 0.128 0.171
-[0.138] [0.186] [0.218] [0.283]
+Treated×$\times$×Post 0.209∗∗∗ 0.336∗∗∗ 0.103 0.255
+[0.070] [0.125] [0.105] [0.200]
 Event-specific ZIP Code FE Y Y Y Y
@@
 Hurricanes Y Y N N
-Observations 48,076 48,076 32,435 32,435
+Observations 62209 62209 43708 43708
 Table 10: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms.
 The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event.
-The coefficient of 0.525 for evictions and 0.928 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 16.9% and 18.4% increases relative to the pre-event averages, respectively.
-The interaction term Treated×$\times$×Post×$\times$×Storm represents the differential impact of storms.
-For evictions, the coefficient is -0.398, not statistically significant, while for filings, the coefficient is -0.757∗∗, indicating a statistically significant reduction in filings in treated ZIP codes exposed to tropical storms.
+The coefficient of 0.331 for evictions and 0.428 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 13.6% and 11.1% increases relative to the pre-event averages of 2.423 and 3.829, respectively.
+The triple interaction Treated×$\times$×Post×$\times$×Storm is negative and marginally significant (-0.228, significant at the 10%) for Eviction, suggesting that the treatment effect after tropical storms is somewhat lower than after hurricanes.
+In other words, tropical storms appear to have a weaker impact on evictions compared to hurricanes.
+However, for eviction filings, the triple interaction (-0.173) is negative but not statistically significant.
 Dependent Variable Eviction Filing
@@
 Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.525∗∗∗ 0.928∗∗∗
-[0.172] [0.235]
-Treated×$\times$×Post×$\times$×Storm -0.398 -0.757∗∗
-[0.271] [0.360]
+Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
+[0.089] [0.110]
+Treated×$\times$×Post×$\times$×Storm -0.228 -0.173
+[0.137] [0.212]
 Event-specific ZIP Code FE Y Y
 Event-specific Time FE Y Y
-Observations 48,076 48,076
+Observations 62209 62209
 3.2 Payday Loans
@@
 We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance.
-We find that when FEMA assistance is made available, transactions volumes remain stable while default rates decrease comparatively, reflecting both a reduction in supply of credit when FEMA is not available but also an increase in precarity in making repayments.
-In particular, Figure 11 and Figure 12 shows that around weeks 6-10 post treatment we observe a spike in defaults and reduction in transaction volumes (possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 15-day loans).
-This also coincides around when evictions peak for non-FEMA receiving hurricanes in the previous section.
-Specifically, in Table 13 we show that the volume of transactions reduces by 8.364 per week and defaults decrease overall by 1.458 per week on average per ZIP code.
-This reflects a 3.8% decrease in transactions and a 14% decrease in defaults per ZIP code per week relative to the baseline period.
-Comparatively, when FEMA rental assistance is not available available, transactions volumes decreased by 27.326 per week, a decline of 12.3%.
-While when FEMA assistance is available, defaults reduced by 1.941 per ZIP code per week relative to the baseline period, a decrease of 18.6% per week per ZIP code relative to the baseline period.
-And as Table 16 show, on the shorter interval of 7 weeks post hurricane event, this difference was even larger, an overall difference of 29.7% in weekly defaults between hurricane events in which FEMA rental assistance was and was not available.
+We find that both transaction volumes and default rates are lower post storm and that when FEMA assistance is made available, transactions volumes and default rates decrease by greater magnitude compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments.
+In particular, Figure 10 we observe a decrease in defaults and reduction in transaction volumes (except a spike around 4,8 and 12 weeks, possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 30-day loans).
+Specifically, in Table 12 we show that the volume of transactions reduces by 8.364 per week and a decrease in defaults by 1.458 per week on average per ZIP code.
+This reflects a 3.8% decrease in transactions per week relative to the baseline period and 14% decrease relative to the baseline period.
+Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 12.597 per week relative to non-FEMA ZIP codes, a decline of 5.7% relative to the pre-event baseline.
+Furthermore, we find a decrease decrease in defaults by 3.318 per week, an approximately 31.8% relative to the pre-event baseline.
+While when ZIP codes receiving FEMA assistance are excluded, transaction volumes reduced by 8.538 per ZIP code per week relative to the baseline period, a decrease of 3.9% per week per ZIP code relative to the baseline period.
+Defaults, on the other hand, show a not statistically significant decrease of 0.433.
 Table 11: Baseline average weekly level per ZIP code in the pre-event period.
@@
 Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
-Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid.
-The dependent variable is the transaction volume.
-Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
-Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
-Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid.
-The dependent variable is the default rates.
-Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
 Table 12: Difference-in-Differences results for transaction volume and default rate outcomes at the ZIP code level for payday loan data.
@@
 Observations 25856 25856 35912 35912
-Table 14: Baseline average weekly level per ZIP code in the pre-event period.
-Variable Control Group Treated Group
-FEMA No FEMA FEMA No FEMA
-Transaction Volume 209.930 283.549 226.306 202.092
-Default 8.286 9.954 11.117 7.738
-Table 15: Difference-in-Differences analysis of transaction volume and default rate outcomes at the ZIP code level, disaggregated by FEMA assistance.
-The analysis compares the impact of the hurricane event on ZIP codes with and without FEMA assistance over two weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
-For the [-5 ; 7] window, ZIP codes with FEMA assistance experienced a small reduction in transaction volume (-2.998), while those without FEMA assistance saw a much larger decrease (-10.675∗∗∗).
-For default rates, FEMA-assisted ZIP codes experienced a slight reduction (-1.837∗∗∗), while those without FEMA saw an increase (1.253∗∗∗).
-For the longer [-5 ; 12] window, FEMA-assisted ZIP codes experienced a smaller reduction in transaction volume (-2.356), while non-FEMA-assisted ZIP codes saw a substantial decrease (-27.237∗∗∗).
-Default rates followed a similar pattern, with a decrease in FEMA-assisted ZIP codes (-1.941∗∗∗) and a small increase in non-FEMA ZIP codes (0.056).
-Weekly Window [-5 ; 7]
+Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level for the treated only.
+The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
+Panel (a) presents transaction estimates for ZIP codes that received FEMA aid and were in the storm path, while panel Panel (b) shows default estimates for ZIP codes that received FEMA aid and were in the storm path.
+Standard errors are clustered at the ZIP code level.
 Dependent Variable Transaction Default
-FEMA No FEMA FEMA No FEMA
-Treated×$\times$×Post -2.998 -10.675∗∗∗ -1.837∗∗∗ 1.253∗∗∗
-[2.750] [1.509] [0.548] [0.327]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Observations 19517 6339 19517 6339
-Weekly Window [-5 ; 12]
+Period Weekly Weekly
+Window [-5 ; 12] [-5 ; 12]
+FEMA×$\times$×Post -12.597 -3.316∗∗∗
+[7.130] [1.112]
+Event-specific ZIP Code FE Y Y
+Event-specific Time FE Y Y
+Observations 7366 7366
+Table 13: Results for transaction volume and default rate outcomes at the ZIP code level for the payday loan data.
+The analysis evaluates the impact of hurricane event across treated groups and interacted with receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
+The coefficient of -12.597 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that received FEMA assistance, corresponding to approximately 5.7% lower transaction volume relative to pre-event levels.
+The coefficient for default rate in the same window is -3.316∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 31.8% lower default rates in treated ZIP codes that received FEMA assistance.
+Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level excluding ZIP codes that received FEMA assistance.
+The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
+Panel (a) presents transaction volume estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance, while panel Panel (b) shows default estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance.
+Standard errors are clustered at the ZIP code level.
 Dependent Variable Transaction Default
-FEMA No FEMA FEMA No FEMA
-Treated×$\times$×Post -2.356 -27.237∗∗∗ -1.941∗∗∗ 0.056
-[2.648] [2.878] [0.531] [0.827]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Observations 27140 8772 27140 8772
-Table 16: Triple Difference-in-Differences (DDD) analysis of transaction volume and default rate outcomes at the ZIP code level, examining the interaction between FEMA assistance and post-event outcomes.
-The table presents results for two weekly windows: [-5; 7] and [-5; 12].
-For both windows, the treatment effect (Treated×$\times$×Post) is negative and statistically significant for transaction volume, indicating a post-event decline in economic activity.
-Conversely, the treatment effect is positive for default rates, suggesting an increase in loan defaults following the event.
-The interaction term (Treated×$\times$×Post×$\times$×FEMA) indicates that FEMA-assisted ZIP codes experienced a significant decrease in default rates in both windows, highlighting the mitigating effect of federal disaster assistance.
-Weekly Window [-5 ; 7] [-5 ; 12]
-Dependent Variable Transaction Default Transaction Default
-Treated×$\times$×Post -10.675∗∗∗ 1.253∗∗∗ -27.237∗∗∗ 0.056
-[1.509] [0.327] [2.877] [0.258]
-Treated×$\times$×Post×$\times$×FEMA 7.677∗∗ -3.090∗∗∗ 24.881∗∗ -1.997∗∗∗
-[3.088] [0.664] [4.142] [0.607]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Observations 25856 25856 35912 35912
+Period Weekly Weekly
+Window [-5 ; 12] [-5 ; 12]
+Treated×$\times$×Post -8.538∗∗∗ -0.433
+[1.727] [0.348]
+Event-specific ZIP Code FE Y Y
+Event-specific Time FE Y Y
+Observations 31686 31686
+Table 14: Difference-in-differences esults for transaction volume and default rate outcomes at the ZIP code level for the payday loan data.
+The analysis evaluates the impact of hurricane event relative to the control group excluding treated ZIP codes receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
+The coefficient of -8.538 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that did not receive FEMA assistance relative to the control group, corresponding to approximately 3.9% lower transaction volume relative to pre-event levels.
+The coefficient for default rate in the same window is -0.433, indicating a non statistically significant reduction in the default rate when excluding treated ZIP codes that received FEMA assistance.
 3.2.2 Hurricanes and Tropical Storms Combined
@@
 We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit.
-Overall, this suggests that major individual level household impacts such as access to housing are sensitive to the availability of emergency liquidity, supporting the need for such programs and the need to understand the availability of other forms of disaster insurance in a world where many locations will be subject to an increasing propensity of climate shock events.
+Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs and the need to better understand their availability in a world subject to an increasing propensity of climate shock events.

### v3

@@
-Keeping in Place After the Storm—Emergency Assistance and Evictions
+
 Abstract
 We offer evidence that federal emergency assistance (FEMA) in the days following natural disasters mitigate evictions in comparison to similar emergency scenarios where FEMA aid is not provided.
-We find an approximate 16.5% increase in overall evictions after hurricane natural disaster events that increases to 19.7% when excluding ZIP codes that receive FEMA rental assistance.
+We find an approximate 10.9% increase in overall evictions after hurricane natural disaster events driven in large part by areas in close proximity of the hurricane path that do not receive FEMA rental assistance.
 Furthermore, we also show that FEMA aid acts as a liquidity buffer to other forms of emergency credit, specifically we find that both transactions volumes and defaults decrease during hurricane events in locations that do receive FEMA aid.
@@
 The loss of structures and in particularly housing is often among the notable and lasting images of in the wake of natural disasters.
-The resulting financial loss and displacement highlight critical questions over both the individual and government role in pre-disaster mitigation and post-disaster response, in particular through their shortcomings.
+The resulting financial loss and displacement highlight the importance of both the individual and government role in pre-disaster mitigation and post-disaster response.
 At the micro level, the loss born by households can be large and has been documented in various forms in the literature.
-Historically, the destruction of housing in such events has often led to rebuilt neighborhoods that differ significantly from their original form and community composition.
-The Boston fire of 1871 (Hornbeck and Keniston, 2017) and the San Francisco 1906 earthquake (Siodla, 2015) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (Boustan et al., 2020) find systematic evidence of out-migration following disaster events in the United States.
+Historically, the destruction of housing in such events has led to rebuilt neighborhoods that differ significantly from their original form and community composition.
+The Boston fire of 1871 (hornbeck2017) and the San Francisco 1906 earthquake (SIODLA201548) provide specific historical moments of migration and reconstruction where entire neighborhoods were rebuilt but also with a notable change in the former population to which (BOUSTAN2020103257) find systematic evidence of out-migration following disaster events in the United States.
 To contrast, in our work, we study instances of assistance in preserving housing access.
@@
 One form of disaster relief has been the need for emergency housing in the wake of damage and destruction to housing property.
-This has further expanded to include rental assistance starting initially under the Stafford Act of 1988 and then further expanded under the Individuals and Households Program in 2002333(Webster, 2022) that aim to recognize disruption in household financial ability due to uninsured or under-insured emergency expenses.
+This has further expanded to include rental assistance starting initially under the Stafford Act of 1988 and then further expanded under the Individuals and Households Program in 2002444(crs_r47015) that aim to recognize disruption in household financial ability due to uninsured or under-insured emergency expenses.
 In the same vein, we study the availability of emergency rental assistance through FEMA in the wake of hurricane disaster events and their impact on rental evictions in the state of Florida.
@@
 The intersection of disasters and eviction risks has also garnered attention.
-(Brennan et al., 2022) explore the heightened vulnerability of renters in disaster-affected areas, showing that disasters exacerbate eviction rates, particularly for vulnerable populations.
-(Raymond et al., 2022) examine the legal and policy dimensions of eviction laws, emphasizing the gaps in tenant protections in states such as Florida.
+(brennan2022perfect) explore the heightened vulnerability of renters in disaster-affected areas, showing that disasters exacerbate eviction rates, particularly for vulnerable populations.
+(raymond2022preventing) examine the legal and policy dimensions of eviction laws, emphasizing the gaps in tenant protections in states such as Florida.
 These findings are particularly relevant to our work, underscoring the cascading effects of disasters on eviction, financial distress, and reliance on high-cost credit.
 Furthermore, the consequences of hurricane events has notably received attention in the literature.
-(Gallagher and Hartley, 2017) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances.
-(Deryugina et al., 2018) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions.
-(Del Valle et al., 2024) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster.
+(gallagher2017household) analyze the financial consequences of hurricane Katrina, highlighting the crucial role of federal aid in stabilizing household finances.
+(deryugina2018economic) corroborate these findings, using tax return data to show long-term income declines and wealth disruptions.
+(del2024household) extend this analysis to hurricane Harvey, examining household decision-making in the aftermath of a disaster.
 They identify shifts in financial behavior, including increased reliance on savings and external credit.
-Similarly, (Collier et al., 2024) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions.
+Similarly, (collier2024credit) demonstrates that FEMA assistance mitigates financial distress, using credit files data they show a reduction in delinquencies and defaults in affected regions.
 Our study complements this result by focusing on evictions and on substitution effects with higher cost credit, particularly how households resort to alternative financial lenders like payday loans when formal aid is absent or insufficient.
 The literature on payday loans has highlighted the dual nature of high-cost credit, as both a financial lifeline and a debt trap.
-We similarly contrast its availability in the context of natural disasters.(Bhutta et al., 2015) describe how payday loans exacerbate financial strain for vulnerable borrowers, while (Morse, 2011) offers a contrasting view, emphasizing their utility during emergencies.
-(Gathergood et al., 2019) shows that payday loans increase financial distress, especially for low-income borrowers, but regulatory interventions can mitigate harmful effects.
-In the same vein, (Dobridge, 2016) finds mixed effects, while access to high-cost credit aids short-term liquidity, it negatively affects long-term financial outcomes.
+We similarly contrast its availability in the context of natural disasters.(bhutta2015payday) describe how payday loans exacerbate financial strain for vulnerable borrowers, while (morse2011payday) offers a contrasting view, emphasizing their utility during emergencies.
+(gathergood2019payday) shows that payday loans increase financial distress, especially for low-income borrowers, but regulatory interventions can mitigate harmful effects.
+In the same vein, (dobridge2016better) finds mixed effects, while access to high-cost credit aids short-term liquidity, it negatively affects long-term financial outcomes.
 Extending this debate, our findings suggest that payday loans act as a substitute for formal disaster relief, especially in regions excluded from FEMA’s coverage.
@@
 Beyond individual outcomes, the broader economic implications of disasters have been extensively analyzed.
-(Bernstein et al., 2019) highlight the long-term economic vulnerabilities of disaster-prone areas, particularly the impact of rising sea levels on property prices.
+(bernstein2019disaster) highlight the long-term economic vulnerabilities of disaster-prone areas, particularly the impact of rising sea levels on property prices.
 The interplay between financial aid and alternative credit sources offers another dimension to understanding recovery dynamics.
-(Collier et al., 2024) analyze demand for disaster recovery loans, illustrating how credit availability shapes household recovery trajectories.
-(Lane, 2024) underscores the importance of guaranteed credit programs in building resilience, particularly for lower-income households.
-(Malmin, 2023) explores the role of federal credit in shaping future wealth trajectories, emphasizing the need for equitable and targeted recovery programs.
+(collier2024credit) analyze demand for disaster recovery loans, illustrating how credit availability shapes household recovery trajectories.
+(lane2024adapting) underscores the importance of guaranteed credit programs in building resilience, particularly for lower-income households.
+(malmin2023access) explores the role of federal credit in shaping future wealth trajectories, emphasizing the need for equitable and targeted recovery programs.
 Finally, recent studies on disaster aid distribution and its long-term effects highlight systemic inequalities.
-(Billings et al., 2022) document the inequities in financial aid distribution during Hurricane Harvey, revealing gaps in support for vulnerable populations.
-(Bufe et al., 2021) explores how financial shocks impact lower-income households, identifying key factors that contribute to financial resilience, such as savings, credit access, and social support networks.
-(Ratnadiwakara and Venugopal, 2020) finds that flood-prone areas increasingly attract lower-income and less creditworthy populations due to declining property values and housing costs.
+(billings2022let) document the inequities in financial aid distribution during Hurricane Harvey, revealing gaps in support for vulnerable populations.
+(bufe2021financial) explores how financial shocks impact lower-income households, identifying key factors that contribute to financial resilience, such as savings, credit access, and social support networks.
+(ratnadiwakara2020areas) finds that flood-prone areas increasingly attract lower-income and less creditworthy populations due to declining property values and housing costs.
 Using Florida as a case study, our research extends prior work by integrating multiple granular datasets to examine the dynamics of financial distress through eviction risks and high-cost credit during the recovery following disasters.
@@
 2.1.1 Eviction Data
-Princeton’s Eviction Lab data provides Census-Tract level information at a weekly level, including eviction, filing, and at-risk outcome variables Desmond et al.
-(2018).
+Princeton’s Eviction Lab data provides Census-Tract level information at a weekly level, including eviction, filing, and at-risk outcome variables evictionlab.
 Covering the period from 2003 to 2017.
@@
 2.1.3 Hurricane Forecast Cone
-The National Hurricane Center (NHC) provides forecast cones for all hurricane and tropical storm events and includes the following characteristics: the forecast date, hurricane category, and the forecast time of impact.
-We use these cones to construct a hurricane time-of-impact and path dataset by examining all the cones for all hurricane and tropical storm events that traverse the state of Florida
+The National Hurricane Center (NHC) provides authoritative information on the timing, location, and progression of tropical storms and hurricanes affecting the United States.
+In our study, this data is critical for reconstructing the geographic footprint of storms that impacted Florida between 2004 and 2018.
 However, the forecast cones are only available as far back as 2008.
-For storm events prior to 2008, we utilize NHC’s data on the storm’s realized path to reconstruct the forecast cone for each storm using the same properties as the forecast cones provided by the NHC in the period after 2008.444https://www.nhc.noaa.gov/aboutcone.shtml provides details on how the forecast cone is constructed.
+For storm events prior to 2008, we relied on NHC forecast cone archives, [^5] https://www.nhc.noaa.gov/aboutcone.shtml provides details on how the forecast cone is constructed.
+which record the storm’s position and projected path at successive forecast times.
+From these records, we extracted the initial forecast points that represent the best estimate of the storm’s current location, thereby tracing the realized path of each hurricane.
 Figure 1 shows an example of a forecast cone.
 The first point represents the actual location of the hurricane, and as we move away from the initial point, the cone covers a larger area with increasing uncertainty, a method of construction chosen by the NHC based on the prediction error over the past five years.
+For storm events prior to 2008, we complemented NHC records with storm track data from NOAA’s Historical Hurricane Tracks database.
+By combining these sources, we were able to build a consistent dataset of storm paths across the full study period.
+This provided the foundation for identifying geographic areas affected by each hurricane and for distinguishing treated from control communities in the empirical analysis.
 Figure 1: 5-Day Track Forecast Cone
@@
 In the following three sections, we will explain how the treated and control groups were constructed and provide a detailed explanation of how the event study and difference-in-differences models were conducted.
-2.2.1 Construction of treatment and control groups
-To measure the impact of hurricanes on households, we construct two groups: treatment and control.
-The treated group consists of ZIP codes (or census tracts) that were within the forecast cone at any point for the entire path of the storm event, from the first prediction to the storm’s dissipation.
-These areas appear in the forecast cone and are likely to experience some direct or/and indirect effects from the hurricane.
-In contrast, our control group includes all ZIP codes (census tracts) that remain completely outside the forecast cone throughout the storm’s progression.
-This ensures that the behavior and outcomes in these areas were more likely to not be influenced by the forecast or the storm itself, allowing us to isolate the impact of the disaster.
-Figures 3 and 3 provide insight into our strategy.
-They show the evolution of four forecast cones for hurricane Hermine in 2016.
-The treated group is the green area, whereas the control group is the intersection between all blue areas, as demonstrated in Figure 4.
-Figure 2: Forecast cones: distinguish between areas inside and outside the cone
-Figure 3: Forecast cones: Evolution of forecast cones during the path of the storm with an example of the treated group (green area)
-Figure 4 provides insight into how we constructed the treatment and control groups.
-Control group (Blue): This group includes ZIP codes (census tracts) that are entirely outside of any given forecast cone, meaning that these areas were not expected to be affected by a hurricane at any point in time.
-By focusing on areas not under the forecast at any point, we construct a control group whose behavior and outcomes are assumed to be unaffected directly by storm warnings or the storm itself.
-Treated Group (Green): This group includes ZIP codes (census tracts) located within the hurricane forecast cone.
-However, only areas that remained in the forecast cone up until the closest forecast cone (3 hours prior to impact) are included.
-That way, we include areas that were eventually either directly in or very close to the realized hurricane path.
-In this setup, we compare two groups that likely have very differential impact.
-We compare the regions that were actually affected (the green area) to the regions (the blue area) that were consistently outside any potential threat of impact.
-The areas in red were excluded as they were previously in the forecast path but did not eventually remain in the realized path of the storm.
-Figure 4: Dynamic 5-Day Forecast Cone
+2.2.1 Construction of treatment and control group
+To assess the impact of hurricanes on households, we classify communities into treatment and control groups based on their geographic proximity to the storm path.
+The treated group consists of ZIP codes whose boundaries intersect the realized hurricane track, extended by a buffer.
+This buffer reflects the forecast uncertainty at the initial advisory position, as defined by the National Hurricane Center (NHC) when constructing its cones of uncertainty.
+It captures not only areas directly along the storm’s path but also nearby regions likely to experience significant effects.
+In this way, the treated group represents communities most plausibly exposed to wind, rainfall, and other storm-related disruptions.
+By contrast, the control group includes ZIP codes in Florida that did not intersect the buffered storm path.
+These areas lie outside the spatial footprint of the hurricane and thus provide a counterfactual for how outcomes would have evolved in the absence of storm exposure.
+This distinction ensures that observed differences between the treated and control groups can be more directly attributed to hurricane impacts rather than broader regional dynamics.
+Figure 2 illustrates our spatial treatment assignment strategy using Hurricane Charley (2004) as an example.
+The map shows how ZIP Code Tabulation Areas (ZCTAs) intersecting the storm’s path are classified as treated, while those outside the exposure zone serve as controls.
+Figure 2: Treatment and Control ZIP Codes for Hurricane Charley (2004).
+The figure shows the hurricane’s observed track (black points), the buffer zone used to define exposure (light blue), and ZIP Codes classified as treated (red) or control (blue).
 2.2.2 Event Study Framework
 To construct the data for the event study, we treat each hurricane and tropical storm as a separate event study.
-We merge the constructed treated and control group dataset of storm affected geographies with the evictions dataset at the census tract level for six months before and six months after the event.
-For the treated group, we use the exact date of the storm path for each census tract and merge with the observed week in the eviction dataset.
-In the control group, since the geographies were constructed to not intersect with the storm forecast, we assign a pseudo-date, in particular, we use the median storm date.
+We merge the constructed treated and control group dataset of storm affected geographies with the evictions dataset at the ZIP code level for six months before and six months after the event.
+For the treated group, we use the exact date of the storm path for each ZIP code and merge with the observed week in the eviction dataset.
+For the control group, whose geographies were constructed to remain outside the storm’s path, we assign a pseudo-date: when the storm event occurred on a single day, we assign that same day to all control ZIP codes, whereas for multi-day events, we randomly distribute the storm dates across control ZIP codes (e.g., if the storm lasted two days, some control ZIP codes are assigned the first day while others are assigned the second).
 Since our evictions dataset is at the weekly level, we calculate the relative event date in weeks and also in months to the storm date observed.
@@
 Together, these terms ensure the model effectively isolates the impact of the event while accounting for unobserved heterogeneity.
-As in (Cengiz et al., 2019) and Dube and Lindner (2024), this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously.
-This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker et al., 2022).
+As in (cengiz) and dube2024minimum, this event study regression aligns events based on event time rather than calendar time, making it comparable to a scenario where all events occur simultaneously.
+This approach aims to correct well documented challenges in the standard difference-in-differences approach (Baker).
 To properly isolate both treated and control groups and to prevent overlap, we only consider a ZIP code as part of the treated group when it is the first event affecting each ZIP code during the event window.
@@
 3.1.1 Hurricanes
-Using our evictions dataset we find that ZIP codes that experienced hurricanes saw a monthly average increase of evictions of 0.525 per month per 1000 renters [^5] We normalize the raw number of evictions and filings using census data on the number of renters present in a ZIP code to account for heterogeneity in the percentage of renters and in the population per ZIP code.
-Table 4.
-This represents a monthly increase of 16% evictions per month compared to the baseline.
+Using our evictions dataset we find that ZIP codes that experienced hurricanes saw a maximum monthly average increase of evictions of 0.201 per month per 1000 renters [^6] We normalize the raw number of evictions and filings using census data on the number of renters present in a ZIP code to account for heterogeneity in the percentage of renters and in the population per ZIP code.
+as in Table 4.
+This represents a monthly increase of 10.9% evictions per month compared to the baseline.
 The event study in Figure 3 shows that this effect persists up to 6 months after the initial hurricane.
-The event study also shows a gradual increase in the initial two months up to a peak around months 3-5, reflecting the nature of the eviction process that may lag from the initial missed rental payment date, to the eviction filing, to the finally court approved eviction order.
-We also evaluate the change in the number of filings, in which we find an increase of 0.928 in the number of filings per month per 1000 renters, in the 6 month that follow the hurricane event.
-This represents a 28.7% increase relative to the baseline rate of filings per month.
-Figure 5: The figure provides the estimated coefficients from an event study examining the effect of hurricane exposure on monthly eviction-related outcomes at the ZIP code level.
-The analysis covers a window from five months before to six months after the hurricane event, with the month immediately preceding the event (month -1) serving as the reference period.
-Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates.
-Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
-Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level.
+The event study also shows a increase around months 2-5, reflecting the nature of the eviction process that may lag from the initial missed rental payment date, to the eviction filing, to the finally court approved eviction order.
+We also show in Table 4 that including ZIP codes by broadening the size of the treatment group radius (in nautical miles) results in a decreasing effect size for evictions, as one might expect given the damage from the hurricane event correlates with distance.
+At 100 nautical miles the effect essentially disappears.
+This result is important as we later show that this effect does not hold when separating into FEMA receiving and non-FEMA receiving ZIP codes.
+Figure 3: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction and filing rates at the ZIP code level with a treatment radius of 21 nautical miles.
+The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period.
+Standard errors are clustered at the ZIP code level.
+Figure 4: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction and filing rates at the ZIP code level with a treatment radius of 50 nautical miles.
+The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period.
+Standard errors are clustered at the ZIP code level.
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Eviction
+[…]
+[0.100] [0.097] [0.103] [0.114] [0.114] [0.084]
+Panel B: Filing
+[…]
+[0.132] [0.126] [0.125] [0.137] [0.150] [0.149]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 22,638 21,617 20,383 19,360 16,939 15,894
+Table 4: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level by treatment radius.
 The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
-the coefficient of 0.331 for evictions represents an approximate 16.5% increase relative to the treated group’s pre-event average.
-Similarly, the 0.428 increase in filings corresponds to about a 13.3% rise from the treated group’s pre-event average.
 Standard errors clustered at the ZIP code level are reported in square brackets.
 ***, **, and * indicate statistical significance at the 1%, 5%, and 10% levels, respectively.
-Dependent Variable Eviction Filing
-Period Monthly Monthly
-Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
-[0.089] [0.110]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 18501 18501
-Table 5: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
-Variable Control Group Treated Group
-Eviction Rate 2.403 2.002
-Filing Rate 3.601 3.203
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Evictions
+Control 2.096 2.112 2.189 2.239 2.302 2.201
+Treated 1.956 1.974 1.842 1.782 1.782 1.824
+Panel B: Filings
+Control 3.129 3.164 3.291 3.372 3.515 3.535
+Treated 3.058 3.041 2.811 2.740 2.746 2.693
+Table 5: Baseline averages by group (pre-hurricane period) for each treatment radius. Control and treated ZIP codes are compared across radii of 10, 15, 21, 26, 50, and 100 miles.
+Figure 5: A bin scatter plot of the proportion of ZIP codes at increasing distances from the hurricane path that had recipients of FEMA rental assistance.
+The plot shows a decreasing in the proportion with a notable jump at the 20 nautical mile mark.
 We then investigate the role of FEMA, in particular hurricane events in which FEMA made renter assistance available in impacted ZIP codes.
-Our results indicate a possible bifurcation in the resultant filings and evictions in the period following the hurricane event.
+Eligibility for assistance through FEMA depends on a geography being declared FEMA eligible [^7] FEMA provides a web page for individuals to lookup whether their ZIP code lies in a FEMA disaster declared county https://www.fema.gov/node/how-do-i-know-if-my-area-eligible-assistance?, as well as individual assessments of damage and hardship which correlate with proximity to the path of the hurricane event.
+Figure 6 shows the ZIP codes that received assistance via FEMA and their distance to the hurricane event.
+We observe that the intersection of ZIP code boundaries, storm damage intensity and receiving FEMA assistance correlate and vary in relation to distance.
+Figure 5 shows the proportion of ZIP codes with individuals receiving FEMA assistance and the distance from the path of the hurricane during FEMA declared hurricanes.
+Figure 6: A map of ZIP codes that received FEMA rental assistance in red during Hurricane Charley.
+The orange non-FEMA treated ZIP codes are those that fall within the hurricane cone but did not receive FEMA rental assistance.
+The map displays two cone sizes, 21 nautical miles and 50 nautical miles, illustrating the sensitivity of the proportion of ZIP codes receiving FEMA rental assistance at different distances.
+Building on our baseline eviction and filings results we observe a bifurcation in the resultant filings and evictions in the period following the hurricane event that varies with FEMA assistance.
 Both the event-study and difference-in-differences regression results illuminate the decrease rate of filings and evictions for hurricane events where FEMA assistance is made available.
-In particular, we firstly find a decrease in the monthly average of 0.418 and 0.445, for evictions and filings per month, respectively Table 6 when restricting to the event study just within treated units.
-These represent a 20.9% and 13.9% decrease, respectively, for hurricanes that did not receive FEMA aid.
-Although both results are not statistically significant, we also show a larger estimate for evictions and filings when excluding ZIP codes that included FEMA assistance in the difference-in-differences results relative to the baseline.
-Comparatively, evictions increased by a statistically significant 0.392 when excluding FEMA receiving ZIP codes relative to the control group, a larger coefficient than 0.331 result when including all ZIP codes.
-Similarly, filings increased by a statistically significant 0.502 when excluding FEMA assistance, larger than the 0.428 coefficient when including all ZIP codes.
-Figure 6: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level on the treated.
-The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period.
-Panel (a) presents the interacted monthly eviction estimates for ZIP codes that received FEMA aid versus those that did not for the treated group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that received FEMA aid.
+In particular, from Table 6 we find a maximal bifurcation in the monthly average evictions with an increase of 0.561 in non-FEMA receiving ZIP codes and a 0.541 increase in filings at the 50 nautical mile radius of inclusion of ZIP codes.
+Figure 7: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction rates at the ZIP code level with a treatment radius of 21 nautical miles.
+The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period.
+The treated group is divided based on FEMA rental assistance availability.
 Standard errors are clustered at the ZIP code level.
-Figure 7: The figure provides the estimated coefficients from an event study evaluating the effect of hurricane exposure on monthly eviction and filing rates at the ZIP code level with the entire set of control ZIP codes but excluding the treated ZIP codes that do not receive FEMA assistance.
-The analysis spans a window from five months before to six months after the hurricane event, with the month preceding the event (month -1) serving as the reference period.
-Panel (a) presents the interacted monthly eviction estimates for ZIP codes that do not receive FEMA assistance versus the control group, while Panel (b) shows monthly interacted filing estimates for ZIP codes that do not receive FEMA assistance versus the control group.
+Figure 8: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on monthly eviction rates at the ZIP code level with a treatment radius of 50 nautical miles.
+The event window spans from five months before to six months after landfall, with the month immediately preceding the hurricane (month -1) serving as the reference period.
+The treated group is divided based on FEMA rental assistance availability.
 Standard errors are clustered at the ZIP code level.
-Dependent Variable Eviction Filing
-Period Monthly Monthly
-Window [-5 ; 6] [-5 ; 6]
-FEMA×$\times$×Post -0.418 -0.445
-[0.291] [0.303]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 2729 2729
-Table 6: Results for monthly eviction and filing outcomes at the ZIP code level on the treated group only, separated into ZIP codes that received FEMA assistance and those did not.
-this provides within storm event estimate comparisons for all ZIP codes that were within the storm path.
-The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
-For FEMA receiving ZIP codes, the coefficient of -0.418, for evictions, while not statistically significant, represents an approximate 20.9% decrease relative to the treated group’s pre-event average (2.002), while the -0.445 decrease in filings corresponds to about a 13.9% decrease from the treated group’s baseline (3.203).
-Dependent Variable Eviction Filing
-Period Monthly Monthly
-Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.392∗∗∗ 0.502∗∗∗
-[0.092] [0.116]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 16647 16647
-Table 7: Difference-in-difference results for monthly eviction and filing outcomes at the ZIP code level while excluding treated ZIP codes that received FEMA assistance.
-This provides a comparison with control group ZIP codes that for the same time period were not in the storm path.
-The estimation window spans 5 months before the event to 6 months after the event, capturing both pre-event and post-event outcomes.
-For non-FEMA receiving ZIP codes, the statistically significant coefficient of 0.392, for evictions, is larger than the 0.331 coefficient for all treated ZIP codes, represents an approximate 19.6% increase relative to the treated group’s pre-event average (2.002), while the 0.502 increase in filings, larger than the 0.428 coefficient for all treated ZIP codes, corresponds to about a 15.6% increase from the treated group’s baseline (3.203).
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Eviction
+[…]
+[0.159] [0.156] [0.164] [0.164] [0.218] [0.141]
+[…]
+[0.106] [0.099] [0.100] [0.111] [0.107] [0.088]
+Panel B: Filing
+[…]
+[0.162] [0.174] [0.186] [0.217] [0.247] [0.215]
+[…]
+[0.140] [0.132] [0.125] [0.131] [0.148] [0.146]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 22,638 21,617 20,383 19,360 16,939 15,894
+Table 6: Difference-in-Differences estimates of post-event eviction and filing outcomes by FEMA-designated and non-FEMA-designated areas across treatment radii.
+Standard errors clustered at the ZIP code level are in brackets.
+***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
 3.1.2 Hurricanes and Tropical Storms
-Table 8: Baseline average monthly level per 1000 renters per ZIP code in the pre-event period.
-Variable Control Group Treated Group
-Eviction Rate 2.664 2.423
-Filing Rate 3.950 3.829
-We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings still holds when lower intensity storms are included.
-Although, the comparative magnitudes decreases, we show that this is largely due to tropical storms seeing far lower post-storm eviction rates.
+We expand our baseline results to include tropical storms to show that the increase in observed evictions and filings when lower intensity storms are included no longer holds.
+We view this as a way to gauge the extent that potential damage severity maybe driving our results, suggesting that the sensitivity to the impact of availability for FEMA assistance matters more within event as opposed to across events.
+In other words, the relative damage of being on the peripheral distances and demand for assistance has observable impacts on evictions relative to any location for lower intensity storms that do not have FEMA assistance at all.
 Hence, our result is likely to be robust from other temporal patterns that may have coincided in our analysis with the comparatively fewer hurricane events.
-Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance.
-Figure 8 and Table 9 shows the baseline result even when including the greater number of tropical storm events still holds (note the three times increase in sample size).
-The decrease is largely driven by the inclusion of tropical storms, noting the treatment post-period coefficient drop from 0.371 to 0.128 for evictions when only tropical storms are included, although the decrease itself is not statistically significant.
-Figure 8: The figure provides the estimated coefficients from an event study examining the effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level.
+Adding additional storm events ensures the increased evictions and filings results are likely due to coinciding with intense hurricane events and the additional absence of FEMA assistance for some geographies.
+Figure 9 and Table 7 shows the baseline result when including the greater number of tropical storm events no longer holds (note the three times increase in sample size).
+Whereas separating the two, shows even a possible statistically significant decline in evictions along the path of tropical storms for radii under 26 nautical miles.
+Figure 9: The figure provides the estimated coefficients from an event study examining the combined effect of storm exposure, including hurricanes and tropical storms on monthly eviction-related outcomes at the ZIP code level.
 The analysis covers a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period.
-Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates.
+Panel (a) displays estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates at the 21 nautical mile treatment radius, whereas Panel (c) displays estimates for monthly eviction rates, while Panel (d) shows estimates for monthly filing rates at the 50 nautical mile treatment radius.
 Vertical lines indicate 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
-Figure 9: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on monthly eviction-related outcomes at the ZIP code level.
-The analysis spans a window from five months before to six months after the storm event, with the month immediately preceding the event (month -1) serving as the reference period.
-Panel (a) reports estimates for monthly eviction rates, while Panel (b) shows estimates for monthly filing rates.
-Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-The results suggest no statistically significant change in either eviction or filing rates following a tropical storm event.
-Table 9: Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms.
-The analysis includes two specifications: one with all storms and another excluding hurricane events.
-The estimation window spans from 5 months before to 6 months after the event.
-In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant increase of 0.209 evictions and 0.336 filings in treated ZIP codes.
-These correspond to approximately 8.6% and 8.7% increases relative to the treated group’s pre-event averages of 2.423 and 3.829, respectively.
-When hurricanes are excluded, the estimated effects are smaller and not statistically significant.
-Dependent Variable Eviction Filing Eviction Filing
-Period Monthly Monthly Monthly Monthly
-Window [-5 ; 6] [-5 ; 6] [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.209∗∗∗ 0.336∗∗∗ 0.103 0.255
-[0.070] [0.125] [0.105] [0.200]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Hurricanes Y Y N N
-Observations 62209 62209 43708 43708
-Table 10: Triple Difference-in-Differences results for monthly eviction and filing outcomes at the ZIP code level, combining both hurricanes and tropical storms.
-The analysis evaluates the differential effects of the tropical storm event across treated and control groups, with an estimation window spanning from 5 months before to 6 months after the event.
-The coefficient of 0.331 for evictions and 0.428 for filings in the Treated×$\times$×Post interaction term suggests a significant increase in both outcomes in the treated ZIP codes, corresponding to approximately 13.6% and 11.1% increases relative to the pre-event averages of 2.423 and 3.829, respectively.
-The triple interaction Treated×$\times$×Post×$\times$×Storm is negative and marginally significant (-0.228, significant at the 10%) for Eviction, suggesting that the treatment effect after tropical storms is somewhat lower than after hurricanes.
-In other words, tropical storms appear to have a weaker impact on evictions compared to hurricanes.
-However, for eviction filings, the triple interaction (-0.173) is negative but not statistically significant.
-Dependent Variable Eviction Filing
-Period Monthly Monthly
-Window [-5 ; 6] [-5 ; 6]
-Treated×$\times$×Post 0.331∗∗∗ 0.428∗∗∗
-[0.089] [0.110]
-Treated×$\times$×Post×$\times$×Storm -0.228 -0.173
-[0.137] [0.212]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 62209 62209
+Treatment Radius
+10 15 21 26 50 100
+Hurricanes and Tropical Storms
+Panel A: Eviction
+[…]
+[0.061] [0.059] [0.054] [0.057] [0.051] [0.043]
+Panel B: Filing
+[…]
+[0.105] [0.099] [0.090] [0.079] [0.072] [0.058]
+Separate Effects: Hurricane and Tropical Storms
+Panel C: Eviction
+[…]
+[0.100] [0.097] [0.103] [0.114] [0.114] [0.084]
+[…]
+[0.075] [0.071] [0.060] [0.060] [0.052] [0.048]
+Panel D: Filing
+[…]
+[0.132] [0.126] [0.125] [0.137] [0.150] [0.149]
+[…]
+[0.114] [0.110] [0.095] [0.078] [0.069] [0.062]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 80,963 83,704 81,921 80,495 75,703 71,206
+Table 7: Difference-in-Differences estimates of eviction and filing outcomes following hurricanes and tropical storms.
+The top panel reports combined treatment effects for all events, while the bottom panels separate effects for hurricanes and tropical storms.
+Standard errors clustered at the ZIP code level are in brackets.
+***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
 3.2 Payday Loans
@@
 We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance.
-We find that both transaction volumes and default rates are lower post storm and that when FEMA assistance is made available, transactions volumes and default rates decrease by greater magnitude compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments.
-In particular, Figure 10 we observe a decrease in defaults and reduction in transaction volumes (except a spike around 4,8 and 12 weeks, possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 30-day loans).
-Specifically, in Table 12 we show that the volume of transactions reduces by 8.364 per week and a decrease in defaults by 1.458 per week on average per ZIP code.
-This reflects a 3.8% decrease in transactions per week relative to the baseline period and 14% decrease relative to the baseline period.
-Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 12.597 per week relative to non-FEMA ZIP codes, a decline of 5.7% relative to the pre-event baseline.
-Furthermore, we find a decrease decrease in defaults by 3.318 per week, an approximately 31.8% relative to the pre-event baseline.
-While when ZIP codes receiving FEMA assistance are excluded, transaction volumes reduced by 8.538 per ZIP code per week relative to the baseline period, a decrease of 3.9% per week per ZIP code relative to the baseline period.
-Defaults, on the other hand, show a not statistically significant decrease of 0.433.
-Table 11: Baseline average weekly level per ZIP code in the pre-event period.
-Variable Control Group Treated Group
-Transaction Volume 228.833 221.300
-Default 8.714 10.418
-Figure 10: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
+In particular, as the individual rental assistance provided by FEMA is of similar magnitudes to the average loan amount available to households via payday loans [^8] The maximum amount in the state of Florida is \$500 and often households borrow at the maximum.
+We find that both transaction volumes and default rates are lower post hurricane event and that this decrease in usage and decrease in defaults is particularly driven when FEMA assistance is made available.
+Both transactions volumes and defaults decrease by greater magnitudes compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments.
+In particular, Figure 11 we observe a decrease in defaults and reduction in transaction volumes.
+Specifically, in Table 8 we show that the volume of transactions reduce by a maximal amount of 17.609 per week and a decrease in defaults by 2.414 per week on average per ZIP code.
+This reflects a 8.3% decrease in transactions per week relative to the baseline period and 22.6% decrease in defaults relative to the baseline period.
+Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 21.209 per week and that non-FEMA ZIP codes see an increase of 13.609 transactions per week at the most maximal, a decline of 10.4% for FEMA receiving ZIP codes and an increase of 6.3% for non-FEMA receiving ZIP codes relative to the pre-event baseline.
+Furthermore, we find FEMA receiving ZIP codes result in a decrease in defaults by 2.810 per week and that non-FEMA ZIP codes no statistically signficant change across almost all treatment radius.
+This represents approximately a 26.3% decrease in defaults relative to the pre-event baseline for FEMA receiving ZIP codes.
+Figure 10: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level with a treatment radius of 21 nautical miles.
+The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
 Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates.
 Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-Table 12: Difference-in-Differences results for transaction volume and default rate outcomes at the ZIP code level for payday loan data.
-The analysis evaluates the impact of hurricane event across treated and control groups over two different weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
-The coefficient of -4.857∗∗ for transaction volume in the [-5 ; 7] window suggests a significant decrease in transaction volume in the treated ZIP codes, corresponding to approximately 2.2% lower transaction volume relative to pre-event levels.
-The coefficient for default rate in the same window is -1.088∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 10.4% lower default rates in treated ZIP codes.
-The longer window, [-5 ; 12], shows even stronger effects with a coefficient of -8.364∗∗∗ for transaction volume, suggesting a 3.8% reduction in transactions, and -1.458∗∗∗ for default rates, corresponding to a 14% reduction in defaults.
-Weekly Window [-5 ; 7] [-5 ; 12]
-Dependent Variable Transaction Default Transaction Default
-Treated×$\times$×Post -4.857∗∗ -1.088∗∗∗ -8.364∗∗∗ -1.458∗∗∗
-[2.040] [0.403] [2.121] [0.416]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Observations 25856 25856 35912 35912
-Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level for the treated only.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
-Panel (a) presents transaction estimates for ZIP codes that received FEMA aid and were in the storm path, while panel Panel (b) shows default estimates for ZIP codes that received FEMA aid and were in the storm path.
-Standard errors are clustered at the ZIP code level.
-Dependent Variable Transaction Default
-Period Weekly Weekly
-Window [-5 ; 12] [-5 ; 12]
-FEMA×$\times$×Post -12.597 -3.316∗∗∗
-[7.130] [1.112]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 7366 7366
-Table 13: Results for transaction volume and default rate outcomes at the ZIP code level for the payday loan data.
-The analysis evaluates the impact of hurricane event across treated groups and interacted with receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
-The coefficient of -12.597 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that received FEMA assistance, corresponding to approximately 5.7% lower transaction volume relative to pre-event levels.
-The coefficient for default rate in the same window is -3.316∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 31.8% lower default rates in treated ZIP codes that received FEMA assistance.
-Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level excluding ZIP codes that received FEMA assistance.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
-Panel (a) presents transaction volume estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance, while panel Panel (b) shows default estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance.
-Standard errors are clustered at the ZIP code level.
-Dependent Variable Transaction Default
-Period Weekly Weekly
-Window [-5 ; 12] [-5 ; 12]
-Treated×$\times$×Post -8.538∗∗∗ -0.433
-[1.727] [0.348]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 31686 31686
-Table 14: Difference-in-differences esults for transaction volume and default rate outcomes at the ZIP code level for the payday loan data.
-The analysis evaluates the impact of hurricane event relative to the control group excluding treated ZIP codes receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]).
-The coefficient of -8.538 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that did not receive FEMA assistance relative to the control group, corresponding to approximately 3.9% lower transaction volume relative to pre-event levels.
-The coefficient for default rate in the same window is -0.433, indicating a non statistically significant reduction in the default rate when excluding treated ZIP codes that received FEMA assistance.
-3.2.2 Hurricanes and Tropical Storms Combined
-Table 15: Baseline average weekly level per ZIP code in the pre-event period for hurricanes and tropical storms combined
-Variable Control Group Treated Group
-Transaction Volume 267.802 256.220
-Default 8.681 9.537
-Table 16: Baseline average weekly level per ZIP code in the pre-event period for tropical storms
-Variable Control Group Treated Group
-Transaction Volume 286.228 291.139
-Default 8.666 8.655
-Adding lower intensity tropical storms once again does not overturn the results in comparing events that include FEMA assistance.
-We continue to observe a decrease in transaction volume and a comparative decrease in default rates for events that saw FEMA assistance.
-Hence, allowing us to control for possible additional temporal events and for differences possibly due to intensity of non-FEMA hurricane events.
-Table 17 and Table 18 show that overall transaction volumes decrease by 7.468 weekly transactions per ZIP code and that defaults decrease by 0.982 defaults per week on average relative to a statistically non-significant change in defaults when only including tropical storm events.
-Most of the change in decreases in transaction volumes and in preservation of default rates is not attributed to non-FEMA events being possibly due to lower intensity storms relative to FEMA declared hurricane events as tropical storms show similar magnitudes of changes in transaction volumes and non-decrease in default rates to non-FEMA hurricanes as in Table 14.
-Figure 13: The figure provides the estimated coefficients from an event study examining the impact of storm (hurricane and tropical storm) exposure on weekly payday loan-related outcomes at the ZIP code level.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
+Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level with a treatment radius of 50 nautical miles.
+The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
 Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates.
 Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-Figure 14: The figure provides the estimated coefficients from an event study examining the impact of tropical storm exposure on weekly payday loan-related outcomes at the ZIP code level.
-The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
-Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates.
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Transaction Volume
+[…]
+[2.962] [2.497] [2.404] [2.228] [2.021] [2.535]
+Panel B: Default
+[…]
+[0.586] [0.479] [0.441] [0.409] [0.404] [0.586]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 45,526 41,821 38,920 36,613 31,904 29,527
+Table 8: Difference-in-Differences estimates of post-event impacts on payday loan transaction volume and defaults across treatment radii.
+Standard errors clustered at the ZIP code level are in brackets.
+***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Transaction Volume
+Control 251.098 259.797 263.990 271.601 291.747 296.754
+Treated 247.778 226.742 227.980 217.360 204.819 212.644
+Panel B: Default
+Control 9.697 9.821 9.713 9.881 10.185 10.085
+Treated 12.010 11.146 11.243 10.566 10.298 10.678
+Table 9: Baseline averages by group (pre-hurricane period) for each treatment radius. Control and treated ZIP codes are compared across radii of 10, 15, 21, 26, 50, and 100 miles.
+Figure 12: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on weekly payday loan transaction volume at the ZIP code level with a treatment radius of 21 nautical miles.
+The event window spans from five weeks before to twelve weeks after landfall, with the week immediately preceding the hurricane (week -1) serving as the reference period.
+The treated group is divided based on FEMA rental assistance availability.
+Standard errors are clustered at the ZIP code level.
+Figure 13: Estimated coefficients from an event study model evaluating the impact of hurricane exposure on weekly payday loan transaction volume at the ZIP code level with a treatment radius of 21 nautical miles.
+The event window spans from five weeks before to twelve weeks after landfall, with the week immediately preceding the hurricane (week -1) serving as the reference period.
+The treated group is divided based on FEMA rental assistance availability.
+Standard errors are clustered at the ZIP code level.
+Treatment Radius
+10 15 21 26 50 100
+Panel A: Transaction Volume
+[…]
+[4.350] [3.096] [2.927] [3.137] [3.096] [3.353]
+[…]
+[3.412] [3.002] [2.879] [2.525] [2.240] [2.597]
+Panel B: Default
+[…]
+[1.211] [0.809] [0.727] [0.701] [0.632] [0.680]
+[…]
+[0.654] [0.553] [0.509] [0.458] [0.440] [0.608]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 45,526 41,821 38,920 36,613 31,904 29,527
+Table 10: Difference-in-Differences estimates of post-event transaction Volume and default outcomes by FEMA-designated and non-FEMA-designated areas across treatment radii.
+Standard errors clustered at the ZIP code level are in brackets.
+***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
+3.2.2 Hurricanes and Tropical Storms Combined
+Adding lower intensity tropical storms for payday loans show a dampened effect for transaction volumes and defaults, unlike the case with evictions where the effect was non-significant.
+In this case, the results remain significant but the magnitudes decrease with treatment radii, which is in line with the a decrease in potential damage from the storm and the absence of FEMA intervention.
+Figure 14 and Table 11 show that overall transaction volumes decrease by a maximal amount of 10.810 weekly transactions per ZIP code and that defaults decrease by 0.504 defaults per week on average.
+Figure 14: The figure provides the estimated coefficients from an event study examining the impact of combined hurricane and tropical storm exposure on weekly payday loan-related outcomes at the ZIP code level.
+The analysis spans a window from five weeks before to twelve weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period.
+Panel (a) reports estimates for weekly transaction volume, while Panel (b) shows estimates for weekly default rates at 21 nautical miles.
+Panel (c) reports estimates for weekly transaction volume, while Panel (d) shows estimates for weekly default rates at 50 nautical miles.
 Vertical lines represent 95% confidence intervals, calculated using standard errors clustered at the ZIP code level.
-Table 17: Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms.
-The analysis includes two specifications: one with all storm events (hurricanes and tropical storms), and another excluding hurricane events.
-The estimation window spans from 5 weeks before to 12 weeks after the event.
-In the full specification including hurricanes and tropical storms, the coefficient on Treated×$\times$×Post indicates a statistically significant decline of 7.468 transactions and 0.932 defaults in treated ZIP codes.
-These correspond to approximately 2.9% and 9.8% decreases relative to the treated group’s pre-event averages of 256.220 transactions and 9.537 defaults, respectively.
-When hurricanes are excluded, the estimated effects are smaller: a 6.345 decrease in transactions (2.2%) and a non-significant 0.273 decrease in defaults (3.2%).
-Weekly Window [-5 ; 12] [-5 ; 12]
-Dependent Variable Transaction Default Transaction Default
-Treated×$\times$×Post -7.468∗∗∗ -0.932∗∗∗ -6.345∗∗∗ -0.273
-[1.316] [0.234] [1.847] [0.200]
-Event-specific ZIP Code FE Y Y Y Y
-Event-specific Time FE Y Y Y Y
-Hurricanes Y Y N N
-Observations 103,291 103,291 67,379 67,379
-${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
-Table 18: Triple Difference-in-Differences results for weekly payday loan transaction volume and default rate outcomes at the ZIP code level, combining both hurricanes and tropical storms.
-The analysis investigates heterogeneity in the treatment effect by interacting the treatment indicator with a binary storm type indicator (1 = tropical storm, 0 = hurricane).
-The estimation window spans from 5 weeks before to 12 weeks after the event.
-The coefficient on Treated×$\times$×Post indicates a statistically significant decrease of 8.364 transactions and 1.458 defaults in treated ZIP codes overall.
-The interaction term Treated×$\times$×Post×$\times$×Storm captures differential effects by storm type: default rates rise after tropical storms relative to hurricanes, potentially reflecting the role of FEMA assistance, which is less commonly triggered by tropical storms.
-Weekly Window [-5 ; 12]
-Dependent Variable Transaction Default
-Treated×$\times$×Post -8.364∗∗∗ -1.458∗∗∗
-[2.040] [0.403]
-Treated×$\times$×Post×$\times$×Storm 2.019 1.185∗∗
-[2.907] [0.463]
-Event-specific ZIP Code FE Y Y
-Event-specific Time FE Y Y
-Observations 103,291 103,291
-${}^{*}p<0.1$, ${}^{**}p<0.05$, ${}^{***}p<0.01$
+Treatment Radius
+10 15 21 26 50 100
+Hurricanes and Tropical Storms
+Panel A: Transaction Volume
+[…]
+[1.696] [1.366] [1.262] [1.163] [0.991] [1.285]
+Panel B: Default
+[…]
+[0.304] [0.246] [0.214] [0.187] [0.146] [0.183]
+Separate Effects: Hurricane and Tropical Storms
+Panel C: Transaction Volume
+[…]
+[2.962] [2.496] [2.404] [2.228] [2.021] [2.535]
+[…]
+[1.886] [1.721] [1.519] [1.447] [1.251] [1.468]
+Panel D: Default
+[…]
+[0.585] [0.479] [0.441] [0.409] [0.404] [0.586]
+[…]
+[0.171] [0.156] [0.140] [0.133] [0.118] [0.176]
+Event-specific ZIP Code FE Y Y Y Y Y Y
+Event-specific Time FE Y Y Y Y Y Y
+Observations 133,020 128,028 131,732 128,489 121,955 116,037
+Table 11: Difference-in-Differences estimates of payday loan outcomes following hurricanes and tropical storms.
+The top panels report combined treatment effects, while the bottom panels separate effects by hurricane and tropical storm exposure.
+Standard errors clustered at the ZIP code level are in brackets.
+***, **, and * indicate significance at the 1%, 5%, and 10% levels, respectively.
 4 Conclusion
@@
 We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit.
-Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs and the need to better understand their availability in a world subject to an increasing propensity of climate shock events.
+Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs while highlighting the need to better understand administrative cutoffs to access.